### v1

# Shrinking the Generation-Verification Gap with Weak Verifiers

Abstract
Verifiers can improve language model (LM) capabilities by scoring and ranking responses from a pool of generated candidates. Currently, high-quality verifiers are either unscalable (e.g., humans) or limited in utility (e.g., tools like Lean for formal proofs). While LM judges and reward models have become broadly useful as general-purpose verifiers, a significant performance gap remains between them and oracle verifiers (i.e. verifiers with perfect accuracy). To help close this gap, we introduce Weaver, a framework for designing a strong verifier by combining multiple weak, imperfect verifiers. First we find that weighted ensembles of verifiers, which typically require learning from labeled data, significantly outperform unweighted combinations due to differences in verifier accuracies. To reduce the dependency on labeled data, Weaver leverages weak supervision to estimate each verifier’s accuracy and combines their outputs into a unified score that better reflects true response quality. However, directly applying weak supervision algorithms poses several challenges, including inconsistent verifier output formats and handling low-quality verifiers. Weaver addresses these challenges by using dataset statistics to normalize outputs and filter specific verifiers. We study the effectiveness of Weaver in test-time repeated sampling settings, where a model generates multiple candidate responses and selects one from among them. Our evaluations demonstrate that Weaver significantly improves over $Pass@1$—the performance when simply selecting the first candidate response—across several reasoning and math tasks, achieving o3-mini-level accuracy with Llama 3.3 70B Instruct (a much cheaper non-reasoning model) as the generator, and an ensemble of 70B or smaller judge and reward models as the verifiers (87.7% average). This gain mirrors the jump achieved between GPT-4o and o3-mini (69.0% vs. 86.7%), which required extensive finetuning and post-training interventions. To reduce the computational costs of running verifier ensembles for Weaver, we train a compact 400M cross-encoder using Weaver’s combined output scores. This distilled model retains 98.7% of Weaver’s full accuracy while reducing verification compute by up to 99.97%.
1 Introduction
A core challenge in deploying language models (LMs) is verification: determining the quality or correctness of a model’s response. This problem arises across various components of the LM pipeline, including dataset curation, model alignment, and inference-time decision-making. Verification relies on verifiers—functions that score responses. When combined with repeated sampling—generating multiple candidate responses from a LM—a perfect verifier can be used to select a correct candidate response, significantly enhancing model capability on tasks such as math, code, and reasoning (snell2024scalingllmtesttimecompute, ; brown2024largelanguagemonkeysscaling, ; puri2025probabilisticinferenceapproachinferencetime, ). For example, Llama 3.1 8B Instruct can match Llama 3.1 70B Instruct and even GPT-4o performances on MATH500 (hendrycks2021measuring, ) and MiniF2F (zheng2022minif2fcrosssystembenchmarkformal, ) when paired with perfect verifiers for these mathematics tasks. However, without a perfect verifier, a generation-verification gap emerges (song2025mindgapexaminingselfimprovement, ): a LM can generate a correct response, but we fail to identify it.
The generation-verification gap is prevalent across many tasks across mathematics, coding, scientific reasoning, instruction-following, and more. For some of these settings, we have access to oracle verifiers that can perfectly identify correct responses. A prominent example is Lean, a formal theorem prover that can be used for problems such MiniF2F (zheng2022minif2fcrosssystembenchmarkformal, ). However, this is often a limited setup, as not all mathematical proofs can be processed by Lean. Alternatively, humans could judge LM responses but manual evaluation is often expensive, noisy, and difficult to scale (hosking2024humanfeedbackgoldstandard, ; clark2021thatshumangoldevaluating, ; karpinska2021perilsusingmechanicalturk, ). In contrast, LMs prompted as judges (chiang2024chatbot, ) and reward models (lambert2024rewardbenchevaluatingrewardmodels, ; singhi2025solveverifycomputeoptimalproblem, ; liu20251bllmsurpass405b, ) can be applied off-the-shelf to tasks like mathematics, coding, scientific reasoning, instruction-following (hendrycks2021measuring, ; rein2024gpqa, ; jain2024livecodebenchholisticcontaminationfree, ; alpaca_eval, ). However, these weak verifiers produce noisy, inconsistent scores, often exhibit poor calibration, and suffer from high false positive rates (stroebl2024inferencescalingflawslimits, ). We ask: to what extent can we leverage weak verifiers to improve accuracy in the repeated sampling regime?
We explore scaling verification, specifically how to combine multiple weak verifiers to improve response selection for repeated sampling. As new pre-trained models become available, the pool of weak verifiers continues to expand and offer diverse, complementary sources of signal that could improve response selection if they can be aggregated effectively. Recent work has explored scaling verification through techniques such as self-verification or averaging LM judge scores (lifshitz2025multiagentverificationscalingtesttime, ; zhao2025sample, ; chen2025sets, ) although other work has found limitations to scaling test-time compute when utilizing weak verifiers for response selection (stroebl2024inferencescalingflawslimits, ). We observe three key challenges towards ensembling weak verifiers:
Figure 1: Weaver Framework: We propose Weaver, a framework combining multiple weak verifiers to effectively scale repeated sampling without parameter finetuning on ground truth labels (left). Weaver significantly outperforms majority voting and shrinks a model’s generation-verification gap by 14.5%, on average, for GPQA Diamond and other datasets (Table 1) (middle). By distilling Weaver from an ensemble of 70B verifiers to a single 400M cross-encoder, we can preserve 98.2% of the accuracy gains of Weaver while reducing inference compute cost by 99.97% (right).
  1. 1.
    Naively aggregating weak verifiers is insufficient for reliable verification. Weak verifiers such as LM-based judges or reward models produce noisy, biased, and poorly calibrated scores, leading to inconsistent performance. (stroebl2024inferencescalingflawslimits, ; lambert2024rewardbenchevaluatingrewardmodels, ; chiang2024chatbot, ). While using a naive unweighted average of verifier scores is straightforward, it implicitly assumes uniform verifier quality, causing low-quality verifiers to dominate and degrade the overall accuracy (verga2024replacingjudgesjuriesevaluating, ; xu2024perfectblendredefiningrlhf, ; eisenstein2023helping, ). Moreover, while previous work has hypothesized that more sophisticated weighted ensembles should perform better, this claim has not been studied (lifshitz2025multiagentverificationscalingtesttime, ).
  2. 2.
    Effective ensembling with limited labeled data is challenging. More sophisticated ensembling techniques typically learn verifier weights from labeled data, but such data is expensive and difficult to obtain. Weak Supervision (WS), a family of statistical techniques developed for data labeling, offers a potential solution through algorithms that aggregate multiple weak signals—such as crowd-worker annotations and expert-defined heuristics—while only requiring a small amount of labeled data (ratner2016data, ; ratner2019training, ; fu2020fast, ). In traditional WS, practitioners can design and shape each weak signal to ensure sufficient quality (i.e., iteratively tweaking program-based heuristics), and guarantees of WS hinge on a baseline level of quality. Our weak signals, however, are fixed pre-trained language model verifiers, which have wildly varying accuracy—especially when applied to out-of-distribution tasks—and can emit incompatible outputs (logits, binary scores, Likert scores) (lambert2024rewardbenchevaluatingrewardmodels, ) that we cannot easily tweak. Due to these conditions, WS algorithms may not perform well when directly applied to verification.
  3. 3.
    Verification is expensive to deploy at inference. Verification can dominate inference-time costs (singhi2025solveverifycomputeoptimalproblem, ; liu20251bllmsurpass405b, ), since each verifier must process both the problem and its candidate response(s) lightman2023letsverifystepstep , often evaluating intermediate steps lightman2023letsverifystepstep and multiple solution paths snell2024scalingllmtesttimecompute . In fact, achieving gains over unverified generation (i.e. majority voting) can require 10×10$\times$10 × to 128×128$\times$128 × the inference compute per query (singhi2025whentosolve, ; lifshitz2025multiagentverificationscalingtesttime, ; zhao2025sample, ; chen2025sets, ).
In this work, we introduce Weaver, a framework for aggregating weak verifiers without supervised finetuning on ground truth labels (Figure 1). First, we demonstrate that if we have access to a large corpus of labeled training data (e.g., 50,000 query-response pairs), we can learn weighted ensembles that can outperform naive averaging by up to 11.2% points. This is because weighted ensembles take advantage of wide variability in verifier accuracy. However, in many real-world scenarios, we do not have access to such quantities of labeled data. Second, to reduce the dependency on labeled data, we adapt Weak Supervision to the verification setting by addressing challenges around inconsistent outputs and low-accuracy verifiers. Weaver filters out uninformative verifiers, normalizes verifier scores, and builds a latent variable model over these scores and the unknown true labels to estimate the verifier accuracies to be used as weights for the ensemble (ratner2016data, ; hall2003generalized, ).
Empirically, given a repeated sampling budget and a set of verifiers, Weaver improves over repeated sampling with unweighted averaging of verifier scores by 17.1% and with majority voting by 13.5% (Table 1; Figure 3). Compared to an LM’s $Pass@1$, Weaver allows us to improve performance by 17.9% for 8B models and 14.5% for 70B models across reasoning and mathematics tasks (Tables 1 and 20). This mirrors the performance jump from GPT-4o to o3-mini (73.9% vs. 88.2%)—but only via increased sampling at test time rather than parameter tuning or post-training procedures. We also study how Weaver scales along different axes of test-time compute: generation, verifiers, model size, and inference budget (Section 5.2). We find that even as we increase the number of generations, many standard verification baselines (e.g. majority voting) quickly plateau (Figure 3). Naive ensembling saturates more slowly, but its gains are limited by sensitivity to the model choice and the number of verifiers.
Finally, to mitigate the compute costs of calling multiple weak verifiers for each response, we extend Weaver by training a 400M-parameter cross-encoder verifier using Weaver’s selected responses. We demonstrate that using a distilled Weaver cross-encoder as a verifier retains 98.7% of the accuracy gains from the learned verifier ensemble while reducing compute costs by three orders of magnitude – saving 99.97% inference FLOPS while still capturing an effective verification strategy (Section 6). Overall, our findings highlight that more reliable, scalable verification is possible even in the absence of ground-truth labels—paving the way for improved data filtering, model alignment, and inference-time decision-making.
2 Related Work
LM Judges and Reward Models: Both LM judges and reward models are promising approaches for evaluating language model outputs, but their high false positive rates limit their reliability (stroebl2024inferencescalingflawslimits, ). LM judges can evaluate outputs without additional training (liu2023g, ; wang2023chatgpt, ; fu2023gptscore, ), using approaches from simple prompting to chain-of-thought reasoning (liu2023g, ) to specialized fine-tuning (saad2023ares, ; tang2024minicheckefficientfactcheckingllms, ) to multi-LM inference architectures (saad2024archon, ; kalra2025verdictlibraryscalingjudgetime, ). However, they face poor generalization across contexts (es2023ragas, ; saad2023ares, ; ravi2024lynxopensourcehallucination, ) and systematic biases in position and self-preference (chen2024humans, ; pan2024human, ; zheng2023evaluation, ; koo2023cognitive, ). Similarly, while reward models have become central to model alignment (bradley1952rank, ; christiano2017deep, ; skyworkreward2024, ), they struggle with noisy training signals from low inter-annotator agreement (askell2021general, ; ouyang2022training, ; wang2024secretsrlhflargelanguage, ; dubois2024alpacafarm, ) and learned biases favoring attributes like response length (lambert2023alignment, ; singhal2023long, ; dubois2024length, ). Recent work has improved individual verifier reliability through better data collection, chain-of-thought reasoning, and natural language unit tests (wang2023helpsteermultiattributehelpfulnessdataset, ; zhang2024generativeverifiersrewardmodeling, ; saadfalcon2024lmunitfinegrainedevaluationnatural, ), yet fundamental challenges persist (eisenstein2023reward, ; chaudhari2024rlhfdecipheredcriticalanalysis, ). Weaver advances beyond these approaches by combining multiple verification signals with adaptive weighting, thus leveraging the complementary strengths of weak verifiers while suppressing noise and reducing false positives.
Weak Supervision: Weaver builds upon statistical techniques from weak supervision, which emerged as a framework for programmatically generating training labels by aggregating multiple weak sources (ratner2016data, ; ratner2020snorkel, ). While a majority of the work focuses on classification tasks (ratner2019training, ; fu2020fast, ; chen2022shoring, ), recent advances have expanded to handle multi-task settings (shin2021universalizing, ) and structured prediction (vishwakarma2022lifting, ). Weak Supervision has also been applied to LM prompting (arora2022askanythingsimplestrategy, ) and routing (guhasmoothie, ). Weaver applies Weak Supervision to answer verification, treating binary imperfect verification signals (e.g. reward models and LM judges) as weak supervision voters that classify candidate solutions as correct or incorrect. This novel application combines predictions by converting these diverse signals into binary verdicts, enabling Weaver to learn better verification strategies from weak but complementary verifiers.
Verification as another compute axis and aggregation: Recent work has explored verification as a new scaling axis lifshitz2025multiagentverificationscalingtesttime ; liu2025can ; zhao2025sample ; singhi2025whentosolve ; stroebl2024inferencescalingflawslimits ; chen2025sets . However this work limits their analysis to one verifier, and instead scale how many times to verify zhao2025sample . Approaches that do leverage multiple verifiers often rely on substantial amounts of labeled data for aggregation or creating specialized verifiers kirchner2024prover ; lifshitz2025multiagentverificationscalingtesttime . With Weaver, we show that it is possible to combine verifiers without ground truth labels, even when they are not specialized. Other work has focused on combining multiple verifiers for post-training the base model using RLHF wang2024transforming ; eisenstein2023helping ; wang2025improvingmodelalignmentcollective .
3 Preliminaries
First, we define the problem of how to select among repeated samples. We then define verifiers and key evaluation metrics, including the generation-verification gap.
Problem Definition    Let $q\in\mathcal{Q}$ be a input text query, and let $r\in\mathcal{R}\sim\mathcal{M}(q)$ be a corresponding response sampled from language model $\mathcal{M}$ with non-zero temperature. For a given query-response pair $(q,r)$, we define $y:\mathcal{Q}\times\mathcal{R}\rightarrow\{0,1\}$ such that $y(q,r)$ is the correctness label of $r$ for $q$.
We are given an unlabeled test dataset $\mathcal{D}^{\text{test}}=\{(q_{i},\bm{r_{i}})\}_{i=1}^{n}$, where $\bm{r_{i}}=\{r_{ij}\}_{j=1}^{K}$ consists of $K$ repeatedly sampled responses from $\mathcal{M}$ for each $q_{i}$. We also assume access to a small labeled development dataset $\mathcal{D}^{\text{dev}}\subset\mathcal{D}^{\text{test}}$, comprising $1\%$ of the test set (e.g. 5 to 10 query-answer pairs), which is used to estimate global statistics such as the task difficulty probability, $\Pr(y_{ij}=1)$. We do not have access to true labels $y{ij}:=y(q_{i},r_{ij})$ for any $i,j$ in $\mathcal{D}^{\text{test}}\setminus\mathcal{D}^{\text{dev}}$.
For each $(q_{i},\bm{r}_{i})\in\mathcal{D}^{\text{test}}$, our goal is to select a correct response $j^{\star}\in[K]$ that satisfies $y_{ij^{\star}}=1$. We can broadly describe this selection rule using a scoring function $f:\mathcal{Q}\times\mathcal{R}\rightarrow\mathbb{R}$, namely $j^{\star}:=\arg\max_{j}f^{\star}(q_{i},r_{ij})$.
Using verifiers    A verifier, either a reward model or an LM prompted as a judge, can be expressed as a scoring function on query-response pairs $v:\mathcal{Q}\times\mathcal{R}\rightarrow\mathbb{R}$. For reward models, the verifier score is continuous, while for LM judges, the verifier score is typically discrete (for our setup, we use $[0,1]$ and $\{0,1\}$, respectively). We assume that we have access to multiple verifiers $\mathcal{V}=\{v_{1},\dots,v_{m}\}$. We apply each of the $m$ verifiers to each $(q_{i},r_{ij})$, for a total of $nmK$ scores on $\mathcal{D}^{\text{test}}$, with $s_{ijk}:=v_{k}(q_{i},r_{ij})$. We aim to use $\mathcal{V}$ to construct a verification strategy $f$.
Evaluation metrics    The $Pass@1$ metric is the probability that an LM’s first response is correct. $Pass@K$ generalizes this metric and is defined as the probability that there exists a correct response among $K$ generated responses: $Pass@K=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(\exists j\in[K]:y_{ij}=1)$. This metric is independent of the verification strategy, and depends on the choice of $\mathcal{M}$, $K$, and the task dataset. The success rate of a verification strategy $\hat{f}$ is $\frac{1}{n}\sum_{i=1}^{n}y_{i\hat{j}}$, where $\hat{j}=\arg\max_{j\in[k]}\hat{f}(q_{i},r_{ij})$. Success rate is dependent on the verification strategy and bounded by Pass@K, and equality is obtained with oracle verification (i.e., $\hat{f}=f^{\star}$ can always select a correct $j$ as long as it exists).
We define the generation-verification gap as Pass@K - Success Rate. A large positive gap indicates that although correct answers are generated, the verification strategy fails to select them consistently. We aim to close this gap and will use it to evaluate verification strategies.
4 Weaver: A Framework for Weak Verifier Aggregation
In Section 4.1, we demonstrate that naively averaging multiple verifier scores to select responses significantly underperforms weighted ensembles; however, common methods for computing weights require labeled data (schapire2013explaining, ; ying2015decision, ). We introduce Weaver (Section 4.2), a method for weighted aggregation of verifier scores with minimal data that draws inspiration from Weak Supervision. Unlike prior work, Weaver adapts weak supervision to verification by addressing challenges unique to verifier aggregation, such as inconsistent score formats and the presence of low-quality or adversarial verifiers. To our knowledge, this is the first framework to successfully apply weak supervision to ensemble verifier scores for response selection.
4.1 How to aggregate multiple verifiers: weighted vs unweighted ensembles
A straightforward approach for using multiple verifiers is a naive ensemble—selecting the response with the highest average verifier score: $f(q_{i},r_{ij})=\frac{1}{m}\sum_{k=1}^{m}s_{ijk}$. This approach lifshitz2025multiagentverificationscalingtesttime does not consider the relative accuracy of verifiers. However, we observed that there is significant variation in the success rates of individual verifiers—spanning a range of up to 37.5%—suggesting that naive ensembles could be suboptimal (Table 16).
An alternative is to use a weighted ensemble. One approach is to use a labeled dataset to identify and use the top-performing verifier, effectively assigning a weight of 00 to discarded verifiers. Other strategies include using Logistic Regression or a Naive Bayes classifier, where the scoring function $f(q_{i},r_{ij})$ is the probability $\Pr(y_{ij}=1|s_{ij1},\dots,s_{ijm})$. These classifiers are fit using labeled data and can be either modeled as a logistic function or factorized using Bayes’ rule and independence assumptions, respectively.
In Figure 2, we compare a naive ensemble with weighted ensembles for several tasks, using Llama 3.3 70B Instruct to generate responses and using a collection of 33 7B-72B reward models and LM judges as verifiers (Appendix C.1). We see that using a weighted ensemble can achieve up to 11.2 points higher success rate than the naive ensemble. However, all weighted ensembles shown are “oracle” methods: they are computed using $y_{ij}$ for all $i\in[n],j\in[K]$, although in practice these labels are unknown for $\mathcal{D}^{\text{test}}$. In fact, when we instead use $0.01n$ labeled samples, accuracy drops by 20.1% on average (Table 17). This raises the question of how to best construct weighted ensembles with limited labeled data.
Figure 2: Weighted Verifier Ensembles Outperform Naive Verifier Ensembles: By using oracle data to keep the best verifiers (i.e. top-$K$ verifier ensembles) or learn aggregation weights for verifiers (i.e. supervised weighted ensembles), we can improve beyond naive combinations of the verifiers available by 3.6% and 7.8%, on average, respectively.
4.2 Weaver: weighted ensembling of verifier scores with minimal labeled data
We first describe the WS method we use in Weaver to construct a weighted ensemble over binary verifier scores. Because verifiers often produce scores in inconsistent formats and exhibit low accuracies—challenges not typically encountered in traditional WS—we introduce a binarization and verifier discarding strategy in Appendices B.2 and B.3 to discard low-quality verifiers and ensure that only sufficiently reliable binary scores are used as input to the WS method.
4.2.1 Weak Supervision Algorithm
In Weak Supervision, the input is an unlabeled dataset, where each entry has multiple binary “votes” on the true label. Applied to our setting, each entry is a query-response pair, forming a dataset of size $nK$, and verifier scores $s_{ijk}$ are binarized into votes $\bar{s}_{ijk}\in\{0,1\}$ for all $i,j,k$. Our goal is to predict the probability that a response is correct, $\Pr(y_{ij}=1|s_{ij1},\dots,s_{ijm})$ for all $i,j$.
WS model
We can view all $y_{ij}$ across query-response pairs as samples of an unknown random variable $Y$ and each $\bar{s}_{ijk}$ across $i,j$ as samples of a random variable $S_{k}$. WS then defines a latent variable graphical model over the random binary vector $\{Y,S_{1},\dots,S_{m}\}$, where $Y$ is latent while $S_{1},\dots S_{m}$ are observable. While existing WS methods assume various models, one common assumption is that $S_{i}\perp S_{j}|Y$ for each $S_{i},S_{j}$. That is, $S_{i}$ and $S_{j}$ are conditionally independent given $Y$; intuitively, each verifier is assumed to capture independent aspects of the correctness of the response (Figure 22 in Appendix C.4). Under this assumption, we can write the posterior probability of a correct generation as the following, for some given binary verifier scores $\{\bar{s}_{1},\dots,\bar{s}_{m}\}$:
$\Pr(Y=1|S_{1}=\bar{s}_{1},\dots,S_{m}=\bar{s}_{m})=\frac{\prod_{i=1}^{m}\Pr(S_{i}=\bar{s}_{i}|Y=1)\Pr(Y=1)}{\Pr(S_{1}=\bar{s}_{1},\dots,S_{m}=\bar{s}_{m})}.$ (1)
The weighted ensemble score for each query-response pair can thus be written in terms of: 1) $\Pr(S_{1}=\bar{s}_{1},\dots,S_{m}=\bar{s}_{m})$, which can be computed from the data; 2) $\Pr(Y=1)$, which can be estimated from $\mathcal{D}^{\text{dev}}$; and 3) $\Pr(S_{i}=\bar{s}_{i}|Y=1)$, or equivalently $\Pr(S_{i}=1|Y=1)$, which is the verifier’s “accuracy parameter”—this cannot be computed directly since we do not have access to $Y$. Next, we discuss how to estimate these accuracy parameters, $\Pr(S_{i}=1|Y=1)$, without labels.
WS parameter estimation
We outline a parameter estimation technique first introduced in (ratner2020snorkel, ). Due to the assumption that $S_{i}\perp S_{j}|Y$, the following equation holds:
$\Pr($ $S_{i},S_{j})=\Pr(S_{i},S_{j}|Y=1)\Pr(Y=1)+\Pr(S_{i},S_{j}|Y=0)\Pr(Y=0)$
$=\Pr(S_{i}|Y=1)\Pr(S_{j}|Y=1)\Pr(Y=1)+\Pr(S_{i}|Y=0)\Pr(S_{j}|Y=0)\Pr(Y=0).$ (2)
Note that $\Pr(S_{i},S_{j})$ can be computed from the known verifier scores, and $\Pr(Y=1)$ is estimated from $\mathcal{D}^{\text{dev}}$. Then, equation 2 is a quadratic equation over the accuracy parameters. We can write this equation for every pair $S_{i},S_{j}$, and for every pair of values $\{0,1\}^{2}$ they can take. Furthermore, we can write another type of equation over the accuracy parameters:
$\Pr(S_{i}=1)=\Pr(S_{i}=1|Y=1)\Pr(Y=1)+\Pr(S_{i}=1|Y=0)\Pr(Y=0).$ (3)
This is a consistency property that holds regardless of the conditional independence assumption, and we can write this equation for each of the $m$ $S_{i}$’s. Because we know that the accuracy parameters should follow equations 2 and 3, we can construct an objective function that aims to minimize the difference between the left and right hand sides of these equations. We write this efficiently in matrix notation. Let $P\in\mathbb{R}^{2\times 2}$ be a diagonal matrix with diagonal $[\Pr(Y=0)\;\Pr(Y=1)]$. Define $\mu\in\mathbb{R}^{m\times 2}$ to be the matrix of accuracy parameters, and define $O\in\mathbb{R}^{2m\times 2m}$ to be a matrix over the joint probabilities of pairs of $S_{i},S_{j}$; more formally:
$\mu_{2i-1:2i,1:2}={\tiny\begin{bmatrix}\Pr(S_{i}=0|Y=0)&\Pr(S_{i}=0|Y=1)\\ \Pr(S_{i}=1|Y=0)&\Pr(S_{i}=1|Y=1)\end{bmatrix}},\;\;O_{2i-1:2i,2i-1:2i}={\tiny\begin{bmatrix}\Pr(S_{i}=0)&0\\ 0&\Pr(S_{i}=1)\end{bmatrix}}\;\forall i\in[m]$
$O_{2i-1:2i,2j-1:2j}={\tiny\begin{bmatrix}\Pr(S_{i}=0,S_{j}=0)&\Pr(S_{i}=0,S_{j}=1)\\ \Pr(S_{i}=1,S_{j}=0)&\Pr(S_{i}=1,S_{j}=1)\end{bmatrix}}\;\forall i\neq j\in[m]$ (4)
Let off-diag denote the elements of a matrix that lie outside its $2\times 2$ block diagonal. Then, to estimate $\mu$ that satisfies both equations 2 and 3, we have the following objective:
$\text{minimize}_{\mu}\bigl{\|}\,O_{\text{off-diag}}-(\mu\,P\,\mu^{T})_{\text{off-diag}}\bigr{\|}^{2}+\bigl{\|}\,\mathrm{diag}(O)-\mu\,P\,\mathbf{1}^{T}\bigr{\|}^{2}$ (5)
We optimize 5 using gradient descent to estimate the verifier accuracy parameters. These estimates are then used in Eq. 1 to select the response with the highest estimated posterior. To further improve modeling of verifier accuracies, we explore whether partitioning the query distribution by empirical difficulty can yield better weak supervision estimates. As detailed in Appendix B.4, we cluster queries based on the observed ratio of correct to incorrect generations, and fit a separate Weaver model within each difficulty bucket. We provide more details in Sec. B.1.
5 Results
In section 5.1, we provide empirical results on Weaver’s performance compared to other approaches for selecting responses in repeated sampling. In section 5.2, we study how Weaver’s performance scales along several axes: the number of responses, model size, verifier counts, and inference compute.
Datasets, Verifiers, and Baselines    Our reward models range in size from 8B to 72B, are all open-source, and are obtained from RewardBench lambert2024rewardbenchevaluatingrewardmodels , a popular evaluation tool for reward models. We prompt open-source language models from Chatbot Arena (chiang2024chatbot, ) to serve as judges. Unless specified, we use Llama 3.3 70B Instruct to generate responses and use all 33 reward models and judges. We evaluate on MATH500, GPQA Diamond, MMLU College, and MMLU Pro. See Sec. C.1 for more details.
We compare Weaver against verifier-free baselines as well as standard verification strategies. First Sample, also known as Pass@1, only uses the first response and does not scale test-time compute or verification. Majority Voting involves repeated sampling but not verification, picking the most common final answer from the responses (brown2024largelanguagemonkeysscaling, ; snell2024scalingllmtesttimecompute, ; chen2024more, ). We compare against the highest scoring reward model and a naive ensemble of the top-10 reward models on RewardBench. We also evaluate two recently proposed methods that scale verification but do not use different verifier models or weighted ensembles: Self-Verification (zhao2025sample, ) and Multi-Agent Verification (lifshitz2025multiagentverificationscalingtesttime, ). Lastly, we report the oracle Pass@K rate, which establishes an upper bound for the success rate of these verification strategies.
5.1 Weaver Shrinks the Gap with Frontier LMs
In Table 1, we evaluate Weaver along with baseline verification methods, the first sample performance of frontier LMs, and the Pass@100 metric. We use LlaMA 3.3 70B Instruct to generate $K=100$ responses per query. We find that Weaver’s weighted ensembling of multiple verifiers allows us to outperform majority vote by $15.5\%$ and come within $4.2\%$ of the Pass@100 oracle metric. Furthermore, Weaver rivals the performance of frontier reasoning models—coming within $0.5\%$ of OpenAI’s o3-mini (OpenAIo3mini2025, )—even though we use a non-reasoning model for generation.
Table 1: Weaver Outperforms Baseline Verification Methods and Shrinks Gap with Frontier LMs.
Methodology Generations ($K$) Datasets Average
MATH
500
GPQA
Diamond
MMLU
College
MMLU
Pro
Baselines First Sample 1 78.0% 42.9% 82.6% 69.9% 68.4%
Majority Voting 100 83.0% 47.4% 84.1% 74.4% 72.2%
Highest Scoring RM on RewardBench (INF-ORM-Llama3.1-70B, ; lambert2024rewardbenchevaluatingrewardmodels, ) 100 78.2% 49.7% 86.0% 77.0% 72.7%
Naive Ensemble of Top-10 RMs on RewardBench (lambert2024rewardbenchevaluatingrewardmodels, ) 100 75.4% 41.3% 88.1% 71.4% 69.1%
Self-Verification (zhao2025sample, ) 100 78.1% 43.1% 82.0% 69.5% 66.9%
Multi-Agent Verification (lifshitz2025multiagentverificationscalingtesttime, ) 100 81.3% 47.8% 84.1% 72.6% 71.6%
Weaver
100 93.4% 72.1% 94.9% 90.2% 87.7%
Frontier Approaches GPT-4o (openaigpt4, ) 1 77.4% 35.9% 87.1% 75.4% 69.0%
Claude 3.7 Sonnet (Anthropic2025Claude37, ) 1 69.2% 48.0% 86.1% 78.1% 70.4%
Llama 4 Maverick (meta2025llama4, ) 1 87.6% 68.9% 91.1% 81.0% 82.2%
o3-mini (OpenAIo3mini2025, ) 1 94.4% 74.0% 92.2% 86.0% 86.7%
Oracle Verification (Pass@100) 100 98.6% 81.0% 96.0% 92.0% 91.9%
5.2 Weaver Improves Compute-Accuracy Trade-Off for Scaling
By proposing to combine multiple weak verifiers instead of one, we introduce yet another axis for test-time scaling. In this section, we study how well scaling verification with Weaver interacts with common previously studied axes for verification, summarized in Table 2.
Table 2: Scaling Dimensions for Generation and Verification Models
Scaling Dimension Base Model Verifier Type Visuals
Sample Count: More Generations Temperature-based sampling Majority Vote, Weak Verifier, Top-K, Weaver Figure 3
Model Size: Larger Models Llama 8B → 70B RM-8B → RM-70B Table 3
Verifier Count: More Models Llama 8B/70B RMs and LM Judges Figure 4
Inference Compute: More FLOPs for Gen./Ver. Temp-based sampling Weak Verifiers + Weaver Figure 5
(1) Scaling Candidate Generations: we study the performance of verification methods as we increase the number of repeated samples in Fig. 3. Based on prior work bradley1952rank ; chen2021evaluating , as the number of responses increases, we are more likely to see a correct response (i.e. Pass@K increases), and hence more likely to select a correct response given a good verification strategy. However, differences in verification translate into different scaling rates. We evaluate the performance of Weaver and baselines for $K=2^{0}$ to $2^{10}$, comparing to o3-mini and Pass@K as well. Across all tasks, Weaver yields the most substantial gains when scaling the number of generations. Weaver consistently narrows the generation-verification gap with the oracle upper bound (Pass@K) while alternative verification strategies plateau after a few generations. The effect is particularly pronounced on difficult tasks like GPQA. We detail the scaling trends observed in Fig. 3 in Sec. C.3.
Figure 3: Scaling Generations Boosts Performance with Weaver: The generation-verification gap shrinks when increasing $K$ and leveraging Weaver, outperforming alternative verification methods by an average 18.3%.
(2) Scaling Model Sizes:
In Table 3, we study how Weaver applied on smaller models (both verifiers and for generating responses) can allow us to match the performance of larger models, enabling weak-to-strong verification. We consider an 8B setting—using LlaMA 3.1 8B to generate responses along with 8B verifiers—and compare this to a 70B setting (LlaMA 3.3 70B Instruct, 8B-72B verifiers) as well as o3-mini. We see that Weaver applied at the 8B scale comes within $1.6\%$ of the majority vote baseline at the 70B scale, and Weaver at 70B surpasses o3-mini by $1.0\%$, demonstrating a weak-to-strong verification phenomenon. Verifier calibration details are available in Appendix C.5.
Table 3: Weaver Reduces Gap between Model Classes: 8B and 70B, 70B and Frontier LM
Generator Model Verifier Model Aggregation Strategy Datasets Average
MATH
GPQA
Diamond
MMLU
College
MMLU
Pro
Llama 3.1 8B Instruct N/A Majority Vote 69.0% 30.5% 72.7% 56.4% 57.2%
8B and below Weaver 80.0% 47.1% 85.7% 67.2% 70.0%
$\Delta$ w. Weaver +11.0% +16.6% +13.0% +10.2% +12.8%
Llama 3.3 70B Instruct N/A Majority Vote 83.0% 44.9% 84.1% 74.4% 71.6%
72B and below Weaver 93.4% 72.2% 94.9% 90.2% 87.6%
$\Delta$ w. Weaver +10.4% +27.3% +10.8% +15.8% +16.0%
o3-mini N/A First Sample 94.4% 74.0% 92.2% 86.0% 86.7%
(3) Scaling Verifier Count:
Two axes for scaling verification are (1) the number of verifiers used and (2) the number of scores sampled from each verifier. Fig. 4 shows how performance changes as we ensemble 1 to 15 verifiers using both naive averaging and Weaver. Verifiers are greedily added in order of individual accuracy, from highest to lowest. Aggregating more verifiers improves performance by up to $8.5\%$ over the top-1 verifier. As shown in Fig. 4, Weaver consistently outperforms naive ensemble averaging across both Oracle Top-5 Verifiers and Total Verifiers configurations for verifier ensembling, with improvements ranging from +2.4% to +10.1% across all datasets. The performance gains are particularly pronounced on GPQA Diamond (+10.1%) and MMLU Pro (+5.1%), demonstrating Weaver’s effectiveness in aggregating verifier signals through learned weights rather than simple averaging. However, gains diminish as more models are added—reflecting the classic ensemble bias-variance tradeoff: initial improvements stem from variance reduction, while additional verifiers contribute redundant signal due to correlated biases on hard examples abebuchanan2024 . We compare alternative score calibration strategies beyond Weaver’s binary transformation in Appendix C.7, and find that the default binarization yields the strongest downstream selection performance. We also explore scaling the number of scores per verifier—via prompt tuning or temperature variation—in Appendix C.5. While this yields modest improvements, increasing verifier count remains the more effective strategy. That said, both methods are complementary and can be combined for further gains.
Figure 4: Weaver Outperforms Naive Ensemble across Oracle Top-5 Verifiers and Total Verifiers Configurations: Results are shown for Weaver ensembles and naive ensembles of the Oracle Top-5 Verifiers (highest-performing verifiers on dataset selected using ground truth) and Total Verifiers (all available verifiers). Weaver consistently outperforms naive ensemble averaging, with improvements ranging from +2.4% to +10.1%.
[…]
(4) Scaling Test-Time Compute:
We study how performance scales in the total compute used for both verification and repeated generations. Figure 5 shows the relationship between inference-time compute and success rate for different generation-verification systems. For each method, we scale the number of generations exponentially from 1 to 100 and plot the required inference compute for generation and verification together versus the success rate. Note that Fig. 5 differs from Fig. 3, since Majority Voting requires 00 verification inference calls while Weaver requires 30+ calls for the weak verifiers. We find that Weaver achieves the highest maximum success rate; notably, majority voting plateaus at around $2^{2}$ to $2^{3}$ ExaFLOPs per query while Weaver continues scaling until 512 ExaFLOPs. However, the additional compute required for Weaver can be prohibitive. We explore how to reduce this computational burden while retaining Weaver’s performance in the next section.
6 Weaver Distillation: Improving Verification Efficiency at Inference
We explore distillation strategies for fine-tuning a smaller LM as a task-specific verifier. In particular, we train cross-encoders; the input is a concatenated query-response pair, while the output is Weaver’s pseudolabel generated from Weak Supervision, namely $\Pr(y_{ij}=1|s_{ij1},\dots,s_{ijm})$ (see Section 4). For the model, we selected ModernBERT-Large (396M) (modernbert, ). For more details, please see Appendix C.6.
Figure 6: Distilling Weaver into a 400M Cross-Encoder Almost Entirely Captures the Performance of Weaver, Yielding 99.97% Compute Savings. We train/evaluate on an 80:20 split.
Figure 6 shows the performance of Weaver on the Llama-70B generations against the cross-encoder on GPQA Diamond. Across tasks, we find that the distilled cross-encoder is able to capture 98.2% of the performance of Weaver. When running Weaver with all the verifiers, it costs 35.35 exaFLOPs for each query’s set of 100 samples. Running a 400M cross-encoder costs 1.01 exaFLOPs for evaluating 100 samples and reduces compute cost by more than three orders of magnitude, saving 99.97% of the FLOPs originally required for running the 70B verifiers. We also outperform majority voting by 23.2% while only incurring a 0.57% increased inference cost over only generating the responses. We see similar results for additional datasets in Figure 22 (Sec. C.6).
These results suggest that, through distillation, we can capture the combined strengths of the weak verifiers used for Weaver, and deploy generalizable and lightweight cross-encoders that use only a fraction of the parameters used for generation. This reduces our hardware constraints considerably; rather than utilizing an 8-GPU node per 70B verifier (i.e. Nvidia H200s with 80B memory), we only require a single A100 GPU with 32GB of memory for our cross-encoder.
7 Discussion
Several research directions remain to be explored with Weaver:
  1. 1.
    Specialized Verifier Development: Our work highlights the varying effectiveness of different weak verifier categories across task domains. Future research should investigate specialized verifier architectures tailored to specific tasks, such as enhanced mathematical reasoning capabilities for numerical problems (yang2024qwen25mathtechnicalreportmathematical, ) or improved code execution simulation for programming tasks (jain2024livecodebenchholisticcontaminationfree, ; quan2025codeelobenchmarkingcompetitionlevelcode, ).
  2. 2.
    Dataset Distribution: For particularly difficult datasets, such as AIMO 2024, Weaver has trouble selecting a correct answer since there are so few correct responses compared to the other datasets (Table 13; Table 14). By scaling the number of generated responses, we are able to improve performance by increasing the absolute number of correct responses (Figure 3) but improved generation techniques, verifier scoring, and aggregation techniques can help us better close the gap with $Pass@K$ for these harder tasks.
  3. 3.
    Weaver for RLHF: With Weaver generated predictions, we can improve the quality of labels used for RLHF on reasoning and mathematics, improving beyond an individual verifier. Previous work has explored reward model ensemble approaches towards RLHF, minimizing poor performing RMs while maximizing the complementary strengths of accurate RMs wang2024transforming ; eisenstein2023helping . Fine-tuning generation models can further improve the accuracy-compute trade-off of Weaver beyond solely distilling a lightweight verifier (Section 6), improving the positive/negative generation ratio and thus making verification an easier task for Weaver models.
  4. 4.
    Multi-Modal Verification: Extending Weaver to multimodal tasks involving images, audio, or video would broaden its applicability but introduces new challenges in verification across modalities (phan2025humanitysexam, ; wang2020vatexlargescalehighqualitymultilingual, ). Research is needed on how verification signals can be effectively combined across different data types.
8 Conclusion
In this paper, we present Weaver, a framework that addresses the fundamental challenge of scaling test-time compute through effective verification strategies. Our contributions advance the state of knowledge in three key dimensions. First, we establish that weighted aggregation of weak verifiers substantially outperforms both individual verifiers and majority voting across reasoning and mathematics tasks, with weighted aggregation exceeding majority voting by an average of 12.3% across all tasks explored (Table 1; Figure 2). Second, we developed a principled approach for unsupervised estimation of verifier accuracies using weak supervision, enabling effective ensemble weighting without fine-tuning on costly ground-truth annotations. This allows us to close the generation-verification gap by 12.8% for 8B models and 16.0% for 70B models (Table 3). By leveraging Weaver with 70B models, we marginally outperform frontier closed-source models such as OpenAI’s o3-mini (87.7% vs. 86.7%) on average across the tasks explored (Table 1). Third, we improve the accuracy-compute trade-off by distilling Weaver into lightweight 400M-parameter cross-encoders. These distilled models retain 98.2% of Weaver’s performance while reducing inference compute by 99.97% (Section 6). This enables high-throughput, cost-efficient verification without sacrificing accuracy, and demonstrates that scalable verification can be achieved without repeatedly querying large models. These findings suggest that strategically combining and distilling weak verifiers enables scalable, label-efficient, and compute-efficient verification—paving the way for better data filtering, model alignment, and inference-time decision-making without additional training of the base generator model.
9 Acknowledgements
We thank the members of the Hazy Lab, Linderman Lab and Scaling Intelligence Lab for their constructive feedback during the composition of the paper.
In particular, we would like to thank Daniel Biderman, Bradley Brown, Ryan Ehrlich, Sabri Eyuboglu, Anna Goldie, Neel Guha, Simon Guo, Jordan Juravsky, Hermann Kumbong, Jerry Liu, Avanika Narayan, Anne Ouyang, Benjamin Spector, Shayan Talaei, Benjamin Viggiano, and Michael Zhang.
We also thank Marlowe and Together AI for providing compute resources that enabled our experiments.
We gratefully acknowledge the support of NIH under No. U54EB020405 (Mobilize); NSF under Nos. CCF2247015 (Hardware-Aware), CCF1763315 (Beyond Sparsity), CCF1563078 (Volume to Velocity), and 1937301 (RTML); US DEVCOM ARL under Nos. W911NF-23-2-0184 (Long-context) and W911NF-21-2-0251 (Interactive Human-AI Teaming); ONR under No. N000142312633 (Deep Signal Processing); Stanford HAI under No. 247183; Google DeepMind; Google Research; Google Cloud; NXP; Xilinx; LETI-CEA; Intel; IBM; Microsoft; NEC; Toshiba; TSMC; ARM; Hitachi; BASF; Accenture; Ericsson; Qualcomm; Analog Devices; Salesforce; Total; the HAI-GCP Cloud Credits for Research program; the Stanford Data Science Initiative (SDSI); members of the Stanford DAWN project: Meta, Google, and VMWare; and members of the Stanford SEAMS project: IBM and Felicis. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views, policies, or endorsements, either expressed or implied, of NIH, ONR, or the U.S. Government.
Appendix A Table of Contents
  1. 1.
    Weaver Methodology (Appendix B)
    1. (a)
      Discrete Weak Supervision Model with Known Difficulty (Appendix B.1)
    2. (b)
      Adapting Weak Supervision to the Verification Setting (Appendix B.2)
    3. (c)
      Filtering Out Low-Quality Verifiers (Appendix B.2.3)
    4. (d)
      Adaptation Method (Appendix B.3)
    5. (e)
      Clustering by Difficulty to Improve Weaver (Appendix B.4)
  2. 2.
    Experiments: (Appendix C)
    1. (a)
      Models and Datasets (Appendix C.1)
    2. (b)
      Verification Baselines (Appendix C.2)
    3. (c)
      Scaling Trends of Weaver (Appendix C.3)
    4. (d)
      Scaling Candidate Generations (Appendix C.4)
    5. (e)
      Scaling Verifier Count (Appendix C.5)
    6. (f)
      Weaver Distillation (Appendix C.6)
    7. (g)
      Individual Verifier Optimization (Appendix C.7)
  3. 3.
    Miscellaneous (Appendix D)
    1. (a)
      Compute Requirements (Appendix D.1)
Appendix B Weaver Methodology
B.1 Weak Supervision Model
We can construct a data generating model over response correctness $y$ and the binary verifier outputs $\bar{s}$. The model is defined as:
Response Correctness: $y_{ij}\sim\text{Bernoulli}(\pi)\;\forall i\in[n],j\in[K],$
Verifier Score: $\bar{s}_{ijk}\mid y_{ij}\sim\begin{cases}\text{Bernoulli}(w_{k,1}),&\text{if }y_{ij}=1,\\ \text{Bernoulli}(1-w_{k,0}),&\text{if }y_{ij}=0,\end{cases}\;\forall i\in[n],j\in[K],k\in[m]$
where:
  • $\pi$ is the probability that a response is correct.
  • $w_{k,1}$ is the true positive rate (TPR) of verifier $k$, and $w_{k,0}$ is the true negative rate (TNR), which we refer to as the verifier’s accuracy parameters.
Here, each verifier $k$ emits a binary score $\bar{s}_{ijk}\in\{0,1\}$, which is assumed to be a noisy indicator of whether response $y_{ij}$ is correct. The likelihood of the verifier’s binary output $X_{ijk}\in\{0,1\}$ is:
$\Pr(S_{k}=\bar{s}_{ijk}\mid Y=y_{ij})=\begin{cases}w_{k,1},&\text{if }y_{ij}=1\text{ and }\bar{s}_{ijk}=1,\\ 1-w_{k,1},&\text{if }y_{ij}=1\text{ and }\bar{s}_{ijk}=0,\\ w_{k,0},&\text{if }y_{ij}=0\text{ and }\bar{s}_{ijk}=0,\\ 1-w_{k,0},&\text{if }y_{ij}=0\text{ and }\bar{s}_{ijk}=1.\end{cases}$
We are interested in estimating the correctness of a response $y_{ij}\in\{0,1\}$ based on assessments from multiple verifiers $\bar{s}_{ij}=\{\bar{s}_{ij1},\bar{s}_{ij2},\dots,\bar{s}_{ijm}\}$. Applying Bayes’ Rule, we get:
$\Pr(y_{ij}=1\mid S=\bar{s}_{ij})=\frac{\Pr(S=\bar{s}_{ij}\mid y_{ij}=1)\Pr(y_{ij}=1)}{\Pr(\bar{s}_{ij})}$ (6)
where $\Pr(\bar{s}_{ij})=\sum_{y^{\prime}\in\{0,1\}}\Pr(\bar{s}_{ij}\mid y_{ij}=y^{\prime})\Pr(y_{ij}=y^{\prime})$.
Eq. 6 requires evaluating the full conditional likelihood:
$\Pr(S=\bar{s}_{ij}\mid y_{ij})=\Pr(S_{1}=\bar{s}_{ij1},S_{2}=\bar{s}_{ij2},\dots,S_{m}=\bar{s}_{ijm}\mid y_{ij}),$
which is a joint distribution over $m$ binary random variables. Since each verifier $S_{k}\in\{0,1\}$ is binary, then there are $2^{m}$ possible verifier output configurations for $S\in\{0,1\}^{m}$. This results in $12^{m}-$ 1 free parameters per class label to construct the distribution $\Pr(S=\bar{s}_{ij}\mid y_{ij})$.
Conditional Independence Assumption
To avoid this exponential blowup, we can assume that the verifiers provide conditionally independent outputs:
$P(S\mid y)=\prod_{k=1}^{m}P(S_{k}\mid y),$
which reduces the number of parameters from $O(2^{m})$ to $O(m)$ and enables efficient inference, under the assumption that each verifier provides unique information about the correctness of a response.
Then, Eq. 6 simplifies to a Naive Bayes-style estimator:
$\Pr\bigl{(}y_{ij}=1\mid S=\bar{s}_{ij}\bigr{)}\;=\;\frac{\Pr(S_{1}=\bar{s}_{ij1},\dots,S_{m}=\bar{s}_{ijm}|y_{ij}=1)\Pr(y_{ij}=1)}{\Pr(S=\bar{s}_{ij})}$
$\;=\;\frac{\Pr(y_{ij}=1)\,\prod_{k=1}^{m}\,\Pr\bigl{(}S_{k}=\bar{s}_{ijk}\mid y_{ij}=1\bigr{)}}{\sum_{y^{\prime}\in\{0,1\}}\Pr(y_{ij}=y^{\prime})\,\prod_{k=1}^{m}\,\Pr\bigl{(}S_{k}=\bar{s}_{ijk}\mid y_{ij}=y^{\prime}\bigr{)}}$ (7)
The parameters in Eq. 7 include:
  • The prior probability of correctness $\pi=\Pr(y_{ij}=1)$.
  • The verifier-specific conditional likelihoods $P(S_{k}\mid y_{ij})$.
B.1.1 Parameter Estimation
Supervised Setting
When ground-truth labels $y_{ij}$ are available, parameter estimation reduces to computing empirical frequencies. We can estimate the prior as:
$\hat{\pi}=\frac{1}{N}\sum_{i,j}\mathbf{1}\{y_{ij}=1\}$
For each verifier $k$, we could estimate:
$\hat{w}_{k,1}$ $=\frac{\sum_{i,j}\mathbf{1}\{y_{ij}=1\}\cdot\mathbf{1}\{S_{k}=1\}}{\sum_{i,j}\mathbf{1}\{y_{ij}=1\}},$
$\hat{w}_{k,0}$ $=\frac{\sum_{i,j}\mathbf{1}\{y_{ij}=0\}\cdot\mathbf{1}\{S_{k}=0\}}{\sum_{i,j}\mathbf{1}\{y_{ij}=0\}}.$
Weak Supervised Setting
When a few labeled $y_{ij}$ are available, we can use it to estimate $\pi$, but we still need to estimate the verifier accuracy parameters $w_{k,1},w_{k,0}$ to compute $\prod_{k=1}^{m}\Pr(S_{k}=\bar{s}_{ijk}|y_{ij}=1)$. Instead of using labeled data, we estimate accuracy parameters using moment matching. In particular, we match observable second moments of verifier outputs to the model-implied moments under conditional independence assumptions, based on an approach from [30].
Pairwise Statistics.
For each pair of verifiers $k_{1},k_{2}$ and binary outputs $a,b\in\{0,1\}$, we can express the joint probability of their outputs using the marginalization rule and the conditional independence assumption:
$\Pr(S_{k_{1}}=a,S_{k_{2}}=b)$ (8)
$=\Pr(S_{k_{1}}=a|Y=1)\Pr(S_{k_{2}}=a|Y=1)\Pr(Y=1)+\Pr(S_{k_{1}}=b|Y=0)\Pr(S_{k_{2}}=b|Y=0)\Pr(Y=0)$
where the conditional distributions for verifier $k$ are:
$\Pr(S_{k}=a\mid y=1)\;=\;\begin{cases}w_{k,1},&a=1,\\ 1-w_{k,1},&a=0,\end{cases}\quad\Pr(S_{k}=a\mid y=0)\;=\;\begin{cases}1-w_{k,0},&a=1,\\ w_{k,0},&a=0.\end{cases}$
Marginal Statistics.
Similarly, each verifier’s marginal distribution can be written as:
$\Pr(S_{k}=1)=\Pr(S_{k}=1|Y=1)\Pr(Y=1)+\Pr(S_{k}=1|Y=0)\Pr(Y=0)$ (9)
Note that this equation holds true regardless of the conditional independence assumption.
Estimation method
  • Construct the second order moment matrix $O\in\mathbb{R}^{(2m)\times(2m)}$, where:
    $O_{2i-1:2i,2i-1:2i}$ $={\tiny\begin{bmatrix}\Pr(S_{i}=0)&0\\ 0&\Pr(S_{i}=1)\end{bmatrix}}\;\forall i\in[m]$
    $O_{2i-1:2i,2j-1:2j}$ $={\tiny\begin{bmatrix}\Pr(S_{i}=0,S_{j}=0)&\Pr(S_{i}=0,S_{j}=1)\\ \Pr(S_{i}=1,S_{j}=0)&\Pr(S_{i}=1,S_{j}=1)\end{bmatrix}}\;\forall i\neq j\in[m]$ (10)
  • Construct the conditional probability matrix $\mu\in\mathbb{R}^{(2V)\times 2}$, where each row encodes:
    $\mu_{2k+a,\,b}\;=\;\Pr\bigl{(}S_{k}=a\mid y=b\bigr{)}$
    $\mu=\begin{bmatrix}w_{k_{1},0}&1-w_{k_{1},1}\\ 1-w_{k_{1},0}&w_{k_{1},1}\\ \dots&\dots\\ w_{k_{m},0}&1-w_{k_{m},1}\\ 1-w_{k_{m},0}&w_{k_{m},1}\end{bmatrix}$
  • Label prior matrix $P\in\mathbb{R}^{2\times 2}$ is a diagonal matrix:
    $P\;=\;\begin{bmatrix}\Pr(y_{ij}=0)&0\\ 0&\Pr(y_{ij}=1)\end{bmatrix}.$
Then, equation 8 is equivalent to $O=\mu P\mu^{\top}$ on the entries off of the $2\times 2$ block diagonal, and equation 9 is equivalent to $\text{diag}(0)=\mu P\mathbb{1}^{\top}$. Therefore, we optimize the following loss to compute $\mu$.
$\text{minimize}_{\mu}\bigl{\|}\,O_{\text{off-diag}}-(\mu\,P\,\mu^{T})_{\text{off-diag}}\bigr{\|}^{2}+\bigl{\|}\,\mathrm{diag}(O)-\mu\,P\,\mathbf{1}^{T}\bigr{\|}^{2},$
By solving via gradient-descent, we obtain estimates of the verifier accuracy parameters $\{w_{k,1},w_{k,0}\}$.
B.1.2 Inference: computing response correctness probabilities
Once the accuracy parameters $\{w_{k,1},w_{k,0}\}$ are estimated and $P(y_{ij}=1)$ is computed from a small labeled development dataset, we can compute posterior correctness probabilities for each response:
$\Pr(y_{ij}=1|S=\bar{s}_{ij})\propto\Pr(y_{ij}=1)\prod_{k=1}^{m}\Pr(S_{k}=\bar{s}_{ijk}|y_{ij}=1)$
$\Pr(y_{ij}=0|S=\bar{s}_{ij})\propto\Pr(y_{ij}=0)\prod_{k=1}^{m}\Pr(S_{k}=\bar{s}_{ijk}|y_{ij}=0)$
Normalizing these, we have a full posterior $P(y_{ij}=1|S=\bar{s}_{ij})$, which provides a score with which we can select a response for each query.
B.2 Adapting Weak Supervision to the Verification Setting
Given a set of verifiers, we elaborate on the design choices behind the weak supervision model described in Sec. 4. In particular, we describe challenges around normalization, binarization and filtering out low-quality verifiers. In Section B.3, we describe Weaver’s approach to normalizing, binarizing, and filtering verifiers.
B.2.1 Normalization
Verifier outputs often differ substantially in scale, range, and distribution. For instance, some verifiers output unbounded real-valued scores (e.g., log-likelihoods), while others output normalized probabilities or learned regression values. Some standard losses under this framework include ranking losses, binary classification losses and regression losses. Some examples include:
Ranking losses
  • Pairwise logistic loss: $\mathcal{L}(s_{1},s_{0})=\log(1+\exp(-(s_{1}-s_{0})))$
  • Bradley–Terry loss: $\mathcal{L}(s_{1},s_{0})=\log(1+\exp(s_{0}-s_{1}))$
  • Triplet margin loss: $\mathcal{L}(s_{1},s_{0})=\max(0,1-(s_{1}-s_{0}))$
Binary classification losses
  • Logistic loss: $\mathcal{L}(s,y)=\log(1+\exp(-y^{\prime}s)),\quad y^{\prime}=2y-1$
  • Hinge loss: $\mathcal{L}(s,y)=\max(0,1-y^{\prime}s)$
Regression losses
  • Squared loss: $\mathcal{L}(s,y)=(s-y)^{2}$
  • Huber loss: $\mathcal{L}_{\delta}(s,y)=\begin{cases}\frac{1}{2}(s-y)^{2}&\text{if }|s-y|\leq\delta\\ \delta(|s-y|-\frac{1}{2}\delta)&\text{otherwise}\end{cases}$
For regression losses, it is essential that the model output $s$ lies in the same range as the label $y\in[0,1]$.
To combine out of the box verifiers, which may be trained under different constraints, verifier scores must be comparable. We note that standard losses imposes different but related invariance assumptions:
  • Ranking and binary classification losses are invariant to positive affine transformations of the form $s\mapsto\alpha s+\beta$, where $\alpha>0$. This means that the relative ordering or decision boundary between scores is preserved even if scores differ in scale or offset.
  • Regression losses, in contrast, directly penalize deviations between scores and targets, so the absolute scale of scores matters. Because many verifier outputs are meant to approximate probabilities or correctness labels $y\in[0,1]$, it is essential to constrain scores to the same interval.
B.2.2 Binarization
The weak supervision algorithm described in the prior section requires binary verifier outputs. This is naturally suited for judge-style verifiers—such as language models prompted to answer yes/no questions—which output discrete ${0,1}$ labels. However, many verifiers, especially reward models, emit continuous scores and often vary in scale and calibration. This raises a key design question: should we input these scores to the weak supervision model as-is, or should we binarize them?
Using continuous scores retains fine-grained information about the confidence of each verifier. This can improve ranking-based performance metrics such as AUC and may allow the weak supervision model to better resolve disagreements among verifiers. However, it introduces challenges when combining signals across verifiers with inconsistent calibration or scale: a score of 0.8 may have different meanings for different verifiers. As seen in Figure Fig. 7 different verifiers exhibit different score distributions even when evaluating the same set of responses. Some verifiers are sharply bimodal, others skew heavily toward low or high scores, and some produce nearly flat or noisy distributions.
To address this, we evaluate several binarization strategies that convert continuous verifier scores into discrete labels. Figure 8 compares the AUC performance of a logistic regression model trained on verifier outputs across four binarization methods: no binarization (continuous scores), a fixed threshold at 0.5, class balance-based thresholds, and quantile binarization. We observe that while continuous scores can achieve strong AUC when sufficient training data is available, simple binarization strategies—especially those that account for score distribution skew—perform comparably and are more robust under limited supervision.
Figure 9 shows only modest differences across binarization strategies for selection accuracy. We note in highly imbalanced datasets, as in the case GPQA, simple quantile-based binarization performs particularly well, likely because it adjusts for the skewed distribution of scores, i.e. it discards ambiguous mid-range scores and retains only the most confident signals.
Figure 7: Accuracy of verifiers on the GPQA dataset. For each verifier, we compute the fraction of problems for which the top-ranked response (according to its score) is correct. While some verifiers consistently select high-quality answers, others perform near chance or worse, motivating the need to filter out low-quality verifiers before applying weak supervision.
Figure 8: Same conventions as Fig. 9 where the performance reported is the area under the curve of the ROC.
Figure 9: Performance for logistic regression models trained on the verifier outputs under different binarization strategies: (1) None: uses raw continuous scores without binarization, (2) Fixed Threshold: applies a uniform threshold across all data, (3) Class Balance: chooses a threshold per verifier so that the proportion of positive labels matches the true class distribution (4) Quantile: assigns positive labels to only the top  15% of scores, focusing on high-confidence predictions. Results are shown across four datasets (GPQA-Diamond, MATH-500, MMLU-College, MMLU-Pro) and four training fractions $(1\%,5\%,10\%,20\%)$. For each seed, we use a random subset of problems as the training set and report the performance on the remaining problems. Curves report mean and standard deviation over multiple seeds. Performance is shown as a function of the number of generations per problem (x-axis, log scale).
B.2.3 Filtering out Low-Quality Verifiers
Verifiers with low accuracy or extreme marginals (e.g., near-constant outputs) not only degrade ensemble performance but also undermine the stability and identifiability of Weak Supervision algorithms, worsening the estimation error. How do we discard verifiers that have low signal?
  • Skewed marginals: Consider a dataset where $\Pr(y=1)\approx 0.5$ and we have a verifier with $\Pr(S_{k}=1)\approx 0.99$. A skewed verifier with an extreme marginal (e.g., from naive thresholding for binarization) and near-constant outputs adds little information to the ensemble. It primarily increases noise in the objective in Eq. 5 and should thus be discarded. Yet, not at all verifiers that have extreme marginals add little signal; for instance, if instead $\Pr(y=1)\approx 0.99$, a skewed verifier could be highly accurate. Therefore, the definition of a low-quality verifier depends on the distribution of correct responses.
  • Breaking symmetry in the WS objective: a common assumption of Weak Supervision is that a majority of the verifiers have better-than-random accuracy [26]. Otherwise, there is a possibility that the WS algorithm can yield non-unique solutions; the terms in Eq. 5 are the joint probabilities over pairs of verifiers as well as their marginals, which do not uniquely determine if a verifier satisfies $w_{k,1},w_{k,0}>0.5$ or not. Therefore, it is critical to remove as many low-accuracy verifiers as possible to ensure that the estimation procedure converges to a unique solution.
B.3 Adaptation Method
Figure 10: Selection accuracy for each verifier across problems and datasets.
Figure 11: Average accuracy for each verifier across problems and datasets given Llama-70B responses scores.
Figure 12: Inverse covariance matrix for each verifier across datasets given Llama-70B responses scores.
Figure 13: Inverse covariance matrix for each verifier across datasets given Llama-70B responses scores, after binarization
Figure 14: Inverse covariance matrix for each verifier across datasets given Llama-70B responses scores, after binarization and dropping.
We now describe our proposed method for normalization, binarization and filtering of verifiers, after which the Weak Supervision algorithm described in Appendix B.1.
  1. 1.
    Normalization: To make verifier outputs comparable, we apply min-max normalization to each verifier:
    $s^{\prime}=\frac{s-\min(s)}{\max(s)-\min(s)}\in[0,1].$
    This ensures that all scores lie within the same numerical range and preserves relative orderings. For regression-style verifiers, normalization aligns their outputs with the scale of the labels and avoids numerical instability due to unbounded score ranges. Without normalization, aggregation methods may become biased or ill-conditioned due to disproportionate score magnitudes.
  2. 2.
    Binarization: We use a small amount of labeled samples $\mathcal{D}^{\text{dev}}$ (which we already are using to compute $\Pr(y=1)$) to determine a threshold for converting continuous verifier outputs to binary outputs. Fig. 13 illustrates the precision matrix of the verifier scores after binarization. It shows a damping of large off-diagonal dependencies and improved condition numbers. Table 4 shows that with only 5 to 10 labeled queries from benchmark development sets (which $\leq$ 1% of the evaluation set), we can estimate binarization thresholds that bolster performance by averages of 8.4% for the 70B models when compared to binary splitting along the median score of the dataset.
  3. 3.
    Filtering out low-quality verifiers: To mitigate the impact of low-quality verifiers, we prune verifiers with extreme marginal behavior, depending on the class balance. For datasets with estimated class balance between 20% and 80%, we filter out verifiers with positive rates outside this range. If a dataset has fewer than 20% positive samples overall, we remove verifiers that predict positives more than 80% of the time. Conversely, for datasets with more than 80% positives, we drop verifiers that predict positives less than 20% of the time.
    Fig. 14 illustrates the precision matrix of the verifier scores after both binarization and dropping low-signal or redundant verifiers. Compared to Fig. 13, it shows further attenuation of off-diagonal structure. Illustrating that dropping contributes substantially to decorrelating the verifier set, which can improve identifiability and numerical stability for downstream weak supervision. As shown in Table 5, verifier pruning leads to 12.5% performance improvement for the 70B model setting.
Table 4: Ablation of Development Set used for Weaver Class Balance Estimation: As our reward model threshold, we set it to the default of the $0.5$.
Approach Model Size Dev Set Size Benchmarks
MATH500 GPQA Diamond MMLU MMLU Pro Average
Weaver 70B
Naive Threshold
(0.5 Threshold)
88.1% 52.0% 92.4% 83.5% 79.0%
Weaver 70B 1% 90.4% 67.1% 91.1% 87.0% 84.5%
Weaver 70B 5% 92.4% 72.7% 93.5% 90.4% 87.5%
Weaver 70B 20% 92.4% 72.7% 93.5% 90.4% 87.5%
Weaver 70B 100% 92.4% 72.7% 93.5% 90.4% 87.5%
Table 5: Ablation of Verifier Selection Strategies for Weaver: Comparison of different strategies for dropping faulty verifiers by dataset using each verifier’s marginal probability.
Approach Model Size Verifier Selection Benchmarks
MATH500 GPQA Diamond MMLU MMLU Pro Average
Weaver 70B No Dropped Verifiers 90.4% 52.0% 91.1% 84.2% 79.4%
Weaver 70B
Low Marginals Dropped
(Mostly Negative Verifiers)
93.4% 60.6% 91.7% 91.0% 84.2%
Weaver 70B
High Marginals Dropped
(Mostly Positive Verifiers)
83.4% 69.7% 87.9% 78.4% 79.9%
Weaver 70B
Extreme Marginals Dropped
(Mostly Positive or Negative)
90.8% 72.7% 92.4% 85.0% 85.2%
Table 6: Ablation of Adaptive Threshold Dev Set Size for Weaver.
Approach Model Size Adaptive Threshold Dev Set Size Benchmarks
MATH500 GPQA Diamond MMLU MMLU Pro Average
Weaver 70B 0.01 92.4% 72.7% 93.5% 90.4% 87.5%
Weaver 70B 0.05 92.4% 72.7% 93.5% 90.4% 87.5%
Weaver 70B 0.2 92.4% 72.7% 93.5% 90.4% 87.5%
Weaver 70B 1.0 92.4% 72.7% 93.5% 90.4% 87.5%
Table 7: Performance Comparison Between Continuous and Discrete Logistic Regression: For supervised fine-tuning on the verifier scores, the continuous model consistently outperforms the discrete variant across all datasets by avoiding the lossy conversion of floats to binary votes required for the discrete variant.
Continuous vs. Discrete Logistic Regression Performance (%)
Method Dataset
MATH 500 GPQA MMLU College MMLU Pro BBH
Discrete LR 93.1 74.3 87.5 87.1 90.1
Continuous LR 97.2 78.1 90.4 92.0 96.5
Improvement +4.1 +3.8 +2.9 +4.9 +6.4
Table 8: Number of Unique Extracted Answers vs. Positive: Negative Sample Ratio per Query - Correlations for Llama 3.1 Instruct Models
Llama 3.1 8B Instruct
Correlation Metrics
Dataset Metric Type
Pearson Spearman Kendall’s Tau
MATH 500 -0.676 -0.745 -0.565
GPQA -0.312 -0.117 -0.096
MMLU College -0.595 -0.700 -0.591
MMLU Pro -0.590 -0.555 -0.425
BBH -0.365 -0.386 -0.300
Llama 3.1 70B Instruct
Correlation Metrics
Dataset Metric Type
Pearson Spearman Kendall’s Tau
MATH 500 -0.631 -0.842 -0.709
GPQA -0.148 -0.093 -0.089
MMLU College -0.551 -0.862 -0.769
MMLU Pro -0.446 -0.693 -0.585
BBH -0.268 -0.594 -0.474
B.4 Exploration: Clustering by Difficulty to Improve Weaver
Weak verifiers often behave inconsistently across the difficulty spectrum of input queries. For instance, most verifiers may have very high accuracy on easy queries and low accuracy on more difficult queries. Fig. 11 and Fig. 10 illustrate the average and selection accuracy of each verifier across multiple problems and datasets. This confirms that there is significant variation in verifier performance at the per-query level. To capture this heterogeneity, we explore clustering queries by difficulty and fitting one Weaver’s weak supervision model per cluster, independently.
We define query difficulty as the empirical ratio of correct to incorrect generations for each query. Using this as a proxy for problem hardness, we partition each dataset into evenly sized clusters along the difficulty distribution and learn separate weak supervision models for each cluster. This is done in an oracle setting, where difficulty is computed using ground-truth correctness, but no label information is used when training the cluster-specific verifier models.
This approach is adaptive in two senses: (1) we adapt the weak supervision model to the difficulty class of the query, and (2) we adapt the threshold of each reward model independently per cluster to better reflect local verifier behavior. For each cluster, we perform a grid search over reward model thresholds ranging from 0.05 to 0.95 in increments of 0.05 (19 values total), selecting the threshold that maximizes accuracy on a held-out development set. We experiment with clustering the queries into between 1 and 5 difficulty levels per dataset, using the oracle difficulty distribution to divide the queries into equally sized bins. While the clustering in our study uses oracle difficulty, future work could explore unsupervised approximations or semi-supervised approaches using a small labeled subset (e.g., 10%) to estimate difficulty distributions.
In Tables 9 and 10,l, we analyze how difficulty-aware clustering and threshold adaptation affect Weaver’s performance at different model scales.
  • 70B model: At this scale, we find that optimizing a single reward model threshold captures most of the verification signal, with clustering yielding only marginal gains ($\sim$1%). This suggests that verifier behavior is relatively stable across query difficulties at higher model capacities.
  • 8B model: In contrast, the 8B setting exhibits a larger gain (4.8%) from clustering and adaptive thresholding. We attribute this to three key factors:
    1. 1.
      Higher verifier variance: As shown in Table 16, verifier quality fluctuates more across queries in the 8B setting, making cluster-specific models more beneficial.
    2. 2.
      Fewer positive generations: Table 13 shows that the 8B generator produces fewer correct answers overall, increasing difficulty heterogeneity.
    3. 3.
      Larger generation-verification gap: The Pass@1-to-Pass@K gap is more pronounced for 8B (e.g., 49.8% to 99.2% on MATH500), indicating greater room for selection-based improvements.
Table 9 confirms that increasing cluster count improves accuracy for the 8B model but often degrades it for the 70B model. These results suggest that difficulty-aware modeling is especially useful when verifier behavior is unstable and when the generation model produces sparse correct candidates.
Further gains from per-model thresholding. Finally, in Table 11, we introduce a finer-grained tuning strategy where each reward model receives its own threshold, rather than using a single global threshold per cluster. This approach provides modest but consistent improvements for the 70B model (e.g., +0.5% on GPQA Diamond, +0.8% on MMLU Pro), and more substantial boosts for the 8B model across all datasets (+1.6% to +2.8%). These gains highlight the value of adapting verifier aggregation strategies not just by query difficulty but also by verifier-specific behavior, especially at smaller model scales where noise is more pronounced.
Table 9: Performance of Weaver with Different Clusters Counts: Utilizes Llama 3.1 70B Instruct Generations with the verifier threshold optimized prior to clustering. We create the clusters based on query difficulty: the ratio of correct to incorrect generations for each queries. We create cluster in evenly sized chunks from the distribution of each task.
Clusters for Weaver Dataset
Dataset Cluster Count
1 2 3 4 5
MATH 500 93.4 87.6 83.8 82.8 81.2
GPQA 66.4 66.4 66.4 66.4 66.4
MMLU College 94.9 91.7 90.1 89.6 89.8
MMLU Pro 88.4 90.2 87.1 84.6 79.8
Average 85.8 84.0 81.9 80.9 79.3
Table 10: Optimizing Clusters and Adaptive Thresholds for Weaver. We report selection performance across different evaluation strategies using weak supervision, with and without difficulty-based clustering and threshold tuning. Clustering is based on oracle query difficulty. Thresholds are selected via grid search from 0.05 to 0.95 in increments of 0.05.
Performance Across Different Evaluation Methods
with Llama 3.1 70B Instruct
Dataset Clustering / Adaptive Threshold
No Clusters /
0.5 Threshold
Best Found
by Search
Pass@K
MATH 500 93.4% 95.2% 98.6%
GPQA Diamond 72.4% 74.1% 81.0%
MMLU College 94.9% 95.1% 96.0%
MMLU Pro 90.2% 90.2% 92.0%
Average 87.7% 88.7% 91.9%
Performance Across Different Evaluation Methods
with Llama 3.1 8B Instruct
Dataset Clustering / Adaptive Threshold
No Clusters /
0.5 Threshold
Best Found
by Search
Pass@K
MATH 500 80.0% 84.3% 99.2%
GPQA Diamond 47.1% 52.7% 95.2%
MMLU College 85.7% 89.9% 98.5%
MMLU Pro 67.2% 72.3% 96.8%
Average 70.0% 74.8% 97.4%
Table 11: Optimizing Clusters and Per-Model Adaptive Thresholds for Weaver. In this setting, each reward model receives its own optimized threshold (rather than a global threshold per cluster). Thresholds are selected via grid search from 0.05 to 0.95 in steps of 0.05. Clustering is still based on oracle query difficulty. This finer-grained tuning yields small improvements for 70B models and more substantial gains for 8B models, especially where verifier accuracy is highly variable.
Performance Across Evaluation Methods with Llama 3.1 70B Instruct (Per-Model Thresholds)
Dataset Clustering / Adaptive Threshold
No Clusters /
0.5 Threshold
Best Found
by Search
Pass@K
MATH 500 93.4% 95.2% 98.6%
GPQA Diamond 72.4% 74.6% 81.0%
MMLU College 94.9% 95.1% 96.0%
MMLU Pro 90.2% 91.0% 92.0%
Average 87.7% 89.0% 91.9%
Performance Across Evaluation Methods with Llama 3.1 8B Instruct (Per-Model Thresholds)
Dataset Clustering / Adaptive Threshold
No Clusters /
0.5 Threshold
Best Found
by Search
Pass@K
MATH 500 80.0% 86.5% 99.2%
GPQA Diamond 47.1% 55.5% 95.2%
MMLU College 85.7% 91.5% 98.5%
MMLU Pro 67.2% 74.5% 96.8%
Average 70.0% 77.0% 97.4%
Appendix C Experiments
C.1 Models and Datasets
Benchmarks: We evaluate our models with several benchmarks for instruction-following, reasoning, mathematics, and coding: MATH500 [31], GPQA [66], MMLU [31], MMLU Pro [89], and BBH [80]. We provide an overview of each dataset in Table 12. For MMLU, we selected the college-level questions for evaluation: biology, chemistry, physics, mathematics, computer science, and medicine. For MMLU Pro, we take a random sample of 500 queries out of the 12K queries available. For BBH, we take four tasks from the dataset of 6K queries available: Penguins in a Table, Causal Judgement, Logical Deduction (Five Objects), and Tracking Shuffled Objects (Five Objects).
Models: We evaluate candidate generations using a range of weak verifiers—models with imperfect but better-than-random accuracy. Our verification system $\mathcal{V}$ includes two primary classes of weak verifiers: Reward Models and LM Judges.
  • Reward Models: A reward model (RM) is a trained language model that assigns a scalar score to candidate responses based on how well they align with human preferences [44, 77]. Given a query and a candidate response, the RM outputs a value $V_{ij}\in[0,1]$ representing the estimated quality of candidate $j$ according to criteria such as correctness, helpfulness, and safety.
    • Examples of reward models include those from the RewardBench leaderboard [44], such as INF-ORM [54], QRM Gemma [19], and Skywork Reward [49]. We also include process reward models (PRMs), which score the reasoning process itself—emphasizing step-by-step logic and coherence—rather than just the final answer [18, 97].
    • For our study, we selected the top-20 reward models from RewardBench and the top-20 process reward models from Process Reward Bench [77] at both 8B and 70B parameter scales. We exclude any RM or PRM that fails to provide a positive learning signal—i.e., those whose rankings perform no better than random selection on benchmark train sets (Appendix  C.1). The diverse training objectives and datasets used for these reward models introduce systematic biases that affect their verification capabilities [44, 77], with different loss functions—including Bradley-Terry loss for pairwise preferences [5], margin loss for fixed score differences [67], and pairwise ranking loss for relative ordering [7].
    • Previous work has noted that it is nontrivial to combine the outputs of reward models and judges as they provide logits and binary decision rules [82, 94]. Instead, we find that we can normalize all RM scores to the range $[0,1]$ using robust percentiles: the bottom 5th percentile is mapped to 0 and the top 95th percentile to 1. For models that provide multiple scoring dimensions (e.g., ArmoRM [85]), we use only their primary output.
  • LM Judges: An LM judge is a language model used to assess the correctness of a candidate response by generating a binary verdict: $V_{ij}\in\{0,1\}$, where 1 indicates that the response is judged correct. These models typically apply chain-of-thought (CoT) reasoning to arrive at their decisions [93]. Each LM judge takes a query and a response as input and outputs a single binary verdict.
    • We use well-known chat models from ChatBotArena [15] as LM judges, which are known for their general-purpose reasoning capabilities. To ensure consistency and determinism, we use greedy decoding (temperature $T=0$) when generating judgments.
Table 12: Benchmark Overview: Evaluation configurations for AlpacaEval 2.0, Arena-Hard-Auto, AIMO, MATH500, GPQA, MMLU, MMLU Pro, and Big-Bench Hard (BBH).
Benchmark Dataset Size Scoring Type Metric License
MATH500 500 Ground Truth Pass@1 Apache 2.0
GPQA 646 Ground Truth Pass@1 CC BY 4.0
MMLU College 719 Ground Truth Pass@1 MIT
MMLU Pro 500 Ground Truth Pass@1 MIT
Figure 15: Growth of Open-Source RMs and LMs: As more and more RMs and LM judges become available, the need for better selection and utilization strategies for these models at test-time continues to grow.
Table 13: Distribution of Generation Accuracy for Llama 3.1 8B Instruct. Each row shows the fraction of queries falling into deciles of correctness—i.e., the proportion of correct generations out of 100 samples per query. The final column reports the overall correct-to-incorrect (C/I) ratio for each dataset. These distributions highlight variation in query difficulty and motivate our clustering-by-difficulty approach in Appendix B.4.
Llama 3.1 8B Instruct
Distribution of Correct/Incorrect Generations
Dataset Percentage of Total Dataset C/I Ratio
0.0-0.1 0.1-0.2 0.2-0.3 0.3-0.4 0.4-0.5 0.5-0.6 0.6-0.7 0.7-0.8 0.8-0.9 0.9-1.0
AIMO 72.2% 7.8% 3.3% 7.8% 2.2% 3.3% 2.2% 0.0% 1.1% 0.0% 10.3%
MATH 500 12.2% 11.4% 11.6% 8.2% 6.4% 8.0% 8.4% 6.60% 11.2% 15.0% 49.9%
GPQA 28.5% 21.2% 16.1% 7.4% 7.3% 5.3% 3.7% 3.3% 2.9% 2.6% 28.3%
MMLU College 7.8% 7.6% 7.1% 6.5% 6.7% 5.3% 7.2% 6.3% 8.6% 22.9% 64.1%
MMLU Pro 22.4% 9.8% 10.6% 5.4% 6.4% 7.2% 4.0% 7.6% 7.6% 15.2% 46.6%
BBH 3.2% 6.7% 10.3% 8.3% 10.3% 14.8% 11.6% 10.7% 10.4% 12.1% 56.9%
Table 14: Distribution of Generation Accuracy for Llama 3.1 70B Instruct. Each row shows the fraction of queries falling into deciles of correctness—i.e., the proportion of correct generations out of 100 samples per query. The final column reports the overall correct-to-incorrect (C/I) ratio for each dataset. These distributions highlight variation in query difficulty and motivate our clustering-by-difficulty approach in Appendix B.4.
Llama 3.1 70B Instruct
Distribution of Correct/Incorrect Generations
Dataset Percentage of Total Dataset C/I Ratio
0.0-0.1 0.1-0.2 0.2-0.3 0.3-0.4 0.4-0.5 0.5-0.6 0.6-0.7 0.7-0.8 0.8-0.9 0.9-1.0
MATH 500 7.0% 4.2% 3.8% 2.0% 2.2% 3.8% 4.2% 4.0% 7.8% 61.0% 78.0%
GPQA 36.8% 5.6% 5.1% 3.7% 6.2% 4.3% 4.5% 4.8% 8.0% 20.9% 42.9%
MMLU College 8.1% 3.2% 1.8% 2.2% 1.5% 1.7% 1.8% 3.2% 2.4% 74.1% 82.6%
MMLU Pro 16.4% 4.2% 1.8% 3.4% 3.0% 3.0% 3.2% 4.8% 6.0% 54.2% 69.9%
Table 15: Models Tested for Weaver.
Model Source Code
Parameter
Count
License Loss Function
LM Judges Llama-3.1-70B-Instruct Open-Source 70B Llama 3.1 Community Cross-Entropy Loss
Llama-3.1-405B-Instruct Open-Source 405B Llama 3.1 Community Cross-Entropy Loss
Llama-3.3-70B-Instruct Open-Source 70B Llama 3.1 Community Cross-Entropy Loss
Meta-Llama-3.1-405B-Instruct-quantized.w8a16 Open-Source 405B Llama 3.1 Community Cross-Entropy Loss
DeepSeek LLM 67B Chat Open-Source 67B DeepSeek License Cross-Entropy Loss
DeepSeekLlama70B Open-Source 70B DeepSeek License Cross-Entropy Loss
DeepSeekQwen32B Open-Source 32B DeepSeek License Cross-Entropy Loss
DeepSeekLlama8B Open-Source 8B DeepSeek License Cross-Entropy Loss
DeepSeekQwen7B Open-Source 7B DeepSeek License Cross-Entropy Loss
Qwen2 72B Instruct Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
Qwen2.5-72B-Instruct Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
Qwen/Qwen2.5-72B-Instruct Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
QwQ-32B Open-Source 32B Apache 2.0 Cross-Entropy Loss
Qwen1.5 110B Chat Open-Source 110B Tongyi Qianwen Cross-Entropy Loss
Qwen1.5 72B Chat Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
Qwen-2.5-7B-Instruct Open-Source 7B Tongyi Qianwen Cross-Entropy Loss
Qwen-2.5-Math-7B-Instruct Open-Source 7B Tongyi Qianwen Cross-Entropy Loss
Mixtral 8x22B v0.1 Open-Source 176B Apache 2.0 Cross-Entropy Loss
Mixtral-8x22B-Instruct-v0.1 Open-Source 176B Apache 2.0 Cross-Entropy Loss
WizardLM 8x22B Open-Source 176B Apache 2.0 Cross-Entropy Loss
WizardLM-2-8x22B Open-Source 176B Apache 2.0 Cross-Entropy Loss
dbrx-instruct Open-Source 132B Databricks Open Model Cross-Entropy Loss
SkyT1 Open-Source 32B Apache 2.0 Cross-Entropy Loss
RMs (8B and below) GRM-Llama3-8B-rewardmodel-ft Open-Source 8B MIT Pairwise Ranking Loss
GRM-Llama3.2-3B-rewardmodel-ft Open-Source 3B Apache 2.0 Pairwise Ranking Loss
GRM-Gemma2-2B-rewardmodel-ft Open-Source 2B Apache 2.0 Pairwise Ranking Loss
Skywork-Reward-Llama-3.1-8B-v0.2 Open-Source 8B Skywork License Pairwise Ranking Loss
QRM-Llama3.1-8B-v2 Open-Source 8B MIT Quantile Regression Loss
URM-LLaMa-3.1-8B Open-Source 8B Skywork License Uncertainty-Aware Loss
GPM-Llama-3.1-8B Open-Source 8B MIT Pairwise Ranking Loss
Llama-3-OffsetBias-RM-8B Open-Source 8B Llama 3.1 Community Pairwise Ranking Loss
ArmoRM-Llama3-8B-v0.1 Open-Source 8B Llama 3.1 Community Pairwise Ranking Loss
Qwen2.5-Math-PRM-7B Open-Source 7B Tongyi Qianwen Cross-Entropy Loss
EurusPRM-Stage1 Open-Source 7B Apache 2.0 Cross-Entropy Loss
EurusPRM-Stage2 Open-Source 7B Apache 2.0 Cross-Entropy Loss
internlm2-7b-reward Open-Source 7B Apache 2.0 Pairwise Ranking Loss
Decision-Tree-Reward-Llama-3.1-8B Open-Source 8B Skywork License Decision Tree Loss
RMs (27B–72B) Skywork-Reward-Gemma-2-27B-v0.2 Open-Source 27B Skywork License Pairwise Ranking Loss
QRM-Gemma-2-27B Open-Source 27B MIT Quantile Regression Loss
INF-ORM-Llama3.1-70B Open-Source 70B Custom License Binary Cross-Entropy Loss
Qwen2.5-Math-RM-72B Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
Qwen2.5-Math-PRM-72B Open-Source 72B Tongyi Qianwen Cross-Entropy Loss
internlm2-20b-reward Open-Source 20B Apache 2.0 Pairwise Ranking Loss
Decision-Tree-Reward-Gemma-2-27B Open-Source 27B Skywork License Pairwise Ranking Loss
Table 16: Weaver Verifier Accuracies and Score Correlations. We report the range of individual verifier accuracies and the average pairwise Pearson correlation between verifier scores. Each verifier’s outputs are flattened across all query–candidate pairs, and correlations are computed across all $\binom{m}{2}$ verifier pairs. Lower correlation indicates greater diversity in how verifiers score responses, which supports the effectiveness of ensembling under Weaver.
Metric Model Size Benchmarks
MATH500 GPQA MMLU Pro Average
Verifier Accuracy Range 8B 34.2% 40.7% 36.4% 37.1%
Avg. Score Correlation 8B 0.0253 0.0349 0.0312 0.0305
Verifier Accuracy Range 70B 27.4% 29.0% 31.6% 29.3%
Avg. Score Correlation 70B 0.0211 0.0372 0.0240 0.0274
C.2 Verification Baselines
C.2.1 Verifier-Free Approaches
First Sample (Pass@1): This baseline uses only the first generated response without any verification or selection mechanism. It represents the standard approach where models generate a single response and provides a lower bound for performance comparison. This method does not scale test-time compute or employ verification.
Majority Voting: A verifier-free approach that generates multiple candidate responses and selects the most frequent final answer across all responses [6, 12, 76]. This method leverages repeated sampling but does not use verification models to assess response quality. Instead, it relies on the assumption that correct answers will appear more frequently than incorrect ones across multiple generations.
C.2.2 Alternative Verification Strategies
Naive Unweighted Aggregation: We consider three oracle configurations using the top-1, top-5, and top-10 verifiers (ranked by their agreement with ground-truth labels). Across all datasets, these oracle ensembles substantially outperform baselines. On average, the best-performing unweighted ensembles exceed first-sample performance by $20.3\%$ and outperform majority voting by $15.0\%$ (see Figure 2). For more difficult benchmarks such as GPQA and MMLU Pro, the top-5 and top-10 ensembles consistently outperform top-1, suggesting that verifier diversity is especially beneficial on challenging examples. However, these oracle ensembles rely on access to ground truth to rank verifiers, limiting their use in practice and motivating the need for learned, unsupervised weighting.
Naive Bayes: We implement a Naive Bayes classifier that models the probability of response correctness given verifier scores: $P(y_{ij}=1|s_{ij1},...,s_{ijm})=\frac{P(s_{ij1},...,s_{ijm}|y_{ij}=1)P(y_{ij}=1)}{P(s_{ij1},...,s_{ijm})}$. Under the conditional independence assumption, this factorizes as $P(s_{ij1},...,s_{ijm}|y_{ij}=1)=\prod_{k=1}^{m}P(s_{ijk}|y_{ij}=1)$. We estimate the parameters using labeled data from the development set. This approach provides a probabilistic framework for aggregating verifier outputs but requires labeled data for parameter estimation.
Logistic Regression: We train a logistic regression classifier where the input features are the verifier scores $[s_{ij1},...,s_{ijm}]$ and the output is the correctness of each response: $P(y_{ij}=1|\mathbf{s}{ij})=\sigma(\mathbf{w}^{T}\mathbf{s}{ij}+b)$, where $\sigma$ is the sigmoid function. The weights $\mathbf{w}$ and bias $b$ are learned using labeled training data. This supervised approach can capture more complex relationships between verifier outputs than naive averaging but requires substantial labeled data for effective training.
Multi-Agent Verification (MAV) [46]: This approach combines multiple "Aspect Verifiers" (AVs) - off-the-shelf LLMs prompted to verify specific aspects of candidate outputs through binary True/False approvals. Unlike reward models, AVs require no additional training and can be easily combined through voting mechanisms. The MAV framework uses BoN-MAV (Best-of-N with Multi-Agent Verification), which: (1) samples $n$ candidate outputs from a generator LLM, (2) collects binary approvals from multiple aspect verifiers that vary across three dimensions (base LLM, aspect to verify, and verification strategy), and (3) selects the output with the most approvals. In our implementation, we use Llama 3.3 70B Instruct as the judge model rather than Gemini 1.5 Flash/Pro as used in the original paper.
Self-Verification [99]: This method implements a sophisticated sampling-based search approach where models verify their own responses through detailed natural language analysis. The approach goes beyond simple self-critique by using structured verification prompts that: (1) rewrite candidate responses in rigorous mathematical theorem-lemma-proof format, (2) systematically scan for errors through step-by-step analysis, and (3) compare responses to localize potential mistakes. The method leverages two key principles: comparing across responses provides signals about error locations (since models struggle with error recall but can identify errors when given their locations), and different output styles are optimal for different tasks (chain-of-thought for generation, rigorous mathematical format for verification). This approach differs from naive self-verification by using structured, multi-step verification protocols rather than simple correctness judgments.
Table 17: Logistic Regression and Naive Bayes Performances across Datasets and Dev Set Sizes
Dataset Approach Model Size Dev Set Size
0.01 0.05 0.2 0.5 1.0
MATH-500 Logistic Regression 70B 70.5% 74.7% 81.4% 93.1% 97.2%
Naive Bayes 70B 67.4% 78.1% 85.0% 89.2% 92.2%
GPQA Diamond Logistic Regression 70B 55.9% 59.4% 69.8% 71.4% 72.9%
Naive Bayes 70B 47.2% 49.2% 57.6% 62.1% 64.3%
MMLU Pro Logistic Regression 70B 72.1% 81.0% 84.6% 86.0% 92.0%
Naive Bayes 70B 60.2% 73.1% 73.1% 78.6% 78.6%
C.3 Scaling Trends of Weaver
Scaling laws describe how performance metrics such as accuracy, sample efficiency or compute cost change as we scale controllable resources, i.e. the number of trials $K$, model capacity. [37] showed that, for fixed-parameter Transformer language models, the cross-entropy loss decreases as a power-law in both model size and data. This framework has since been extended to explore optimal tradeoffs between model and data scaling [33], as well as inference-time scaling with multiple samples [13, 6].
First, we establish the power law scaling of the Pass@K rate. Assume the $i$-th problem has an unknown “difficulty” $p_{i}\in[0,1]$, the probability that one response is correct. With $K$ independent samples, the chance we get at least one correct response is
$q_{i}(p_{i},K)=1-(1-p_{i})^{K}\approx 1-\exp(-p_{i}\cdot K)\quad\text{for small }p_{i}$
Define the indicator variable:
$X_{i}=\begin{cases}1,&\text{if the i-th query is solved at least once (with probability $q_{i}$)},\\ 0,&\text{otherwise},\end{cases}$
and let the total number of solved problems be $Y=\sum_{i=1}^{N}X_{i}$.
[…]
To model population-level variation in problem difficulty, we assume each problem’s correctness probability $p_{i}$ is drawn from a Beta distribution: $p_{i}\sim\text{Beta}(\alpha,\beta)$. This captures the idea that some problems are easier (high $p_{i}$) while others are harder (low $p_{i}$), with the overall distribution controlled by the shape parameters $\alpha,~{}\beta$. Then, the fraction of problem that can be solved in $K$ attempts follows,
Pass@K $=\mathbb{E}_{p\sim\text{Beta}(\alpha,~{}\beta)}[1-(1-p)^{K}]$
$=1-\mathbb{E}_{p\sim\text{Beta}(\alpha,~{}\beta)}[(1-p)^{K}]=1-\frac{B(\alpha,~{}\beta+K)}{B(\alpha,~{}\beta)}$ (11)
by the definition of the Beta function $B(\cdot,\cdot)$. Taking logarithm:
$\log\text{Pass@K}=\log\left(1-\frac{B(\alpha,~{}\beta+K)}{B(\alpha,~{}\beta)}\right)\approx-\frac{B(\alpha,~{}\beta+K)}{B(\alpha,~{}\beta)}$
by $\log(1-x)\approx-x$ when $x$ is small, which holds for large $K$. Then, expressing the Beta function in terms of the Gamma function leads to:
$\log\text{Pass@K}\approx-\frac{\Gamma(\beta+K)\Gamma(\alpha+\beta)}{\Gamma(\beta)\Gamma(\alpha+\beta+K)}$
For large $K$, we can apply Stirling’s approximation of the Gamma function $\log\Gamma(x)\approx x\log x-x+\frac{1}{2}\log(2\pi)+\frac{1}{2}\log x$:
[…]
$\approx(\beta+K)\log(\beta+K)-(\alpha+\beta+K)\log(\alpha+\beta+K)+\frac{1}{2}\log\left(\frac{\beta+K}{\alpha+\beta+K}\right)$
$\approx(\beta+K)\log K-(\alpha+\beta+K)\log K+\text{const}$
$=-\alpha\log K+\log\zeta$
when we retain the leading term. In turn, the log of the expected coverage follows a power law in $K$, scaling as:
$\log\text{Pass@K}=-\exp\left(-\alpha~{}\log K+\log\zeta\right)=-\zeta K^{-\alpha}$ (12)
Verifier Success Modeling
Now suppose we pass the $K$ candidates through a scoring model ("verifier”) which selects the top-scoring answer. The verification process succeeds if (i) at least one correct answer was generated and (ii) the verifier ranks a correct answer highest.
[…]
Assume the verifier assigns scores such that correct responses are drawn from a score distribution $f_{1}$, and the incorrect responses from a distribution $f_{0}$. Let $s^{(1)}=\{s_{j}:y_{j}=1\}$ and $s^{(0)}=\{s_{j}:y_{j}=0\}$ denote the scores of correct and incorrect responses, respectively. Then a query is successfully verified if:
[…]
Our goal is to compute the probability that the maximum of $c$ i.i.d draws from $f_{1}$ exceeds the maximum of $K-c$ draws from $f_{0}$.
To model the correctness of responses, we assume each query $i$ has a latent correctness probability $p_{i}\sim\text{Beta}(\alpha,\beta)$, reflecting query-specific difficulty. Given $p_{i}$, each of the $K$ responses is sampled independently as:
$y_{ij}\sim\text{Bernoulli}(p_{i}),\quad j=1,\dots,K$
This implies the number of correct responses follows a Binomial distribution:
$C_{i}=\sum_{j=1}^{K}y_{ij}\sim\text{Binomial}(K,p_{i})$
assuming (1) conditional independence of responses given $p_{i}$, (2) identical correctness probabilities within a query, and (3) a fixed number of responses $K$.
Because the correctness probability $p$ varies across queries, the dataset-level Selection@1 curve requires marginalizing over $p$:
[…]
Combined with the need to model max comparisons over verifier scores, it renders the exact calculation of Selection@1 analytically intractable.
To enable tractable, smooth modeling of Selection@1, we introduce the following parametric form:
[…]
  • The coverage term $\exp(-\zeta K^{-\alpha})$ approximates the probability that at least one correct response is generated.
  • The verification term $1-(1-\pi)^{K^{\gamma}}$ approximates the chance that the top-scoring response is correct, given that at least one correct response exists. The parameter $\gamma$ controls whether verifier performance improves sublinearly or superlinearly with $K$. The parameter $\pi$ represents the effective per-response probability that a correct response is successfully selected by the verifier, conditioned on the response being correct and included in the candidate set.
To obtain practical scaling trends, we fit parametric models in Eq. 14 to the empirical averages computed from $5$ independent runs for each value of $K$, across each dataset and verification strategy. Specifically, we use the L-BFGS-B algorithm to optimize a smooth approximation following [32]. To ensure numerical stability and robustness to outliers or heavy-tailed noise in the observed selection accuracies, we minimize the Huber loss between the predicted values and the empirical means. The Huber loss behaves quadratically for small residuals and linearly for large ones, making it less sensitive to outliers than mean squared error (MSE) while maintaining smooth differentiability for gradient-based optimization. It is defined as,
$L_{\delta}(r)=\begin{cases}\frac{1}{2}r^{2}&\text{if }|r|\leq\delta\\ \delta\left(|r|-\frac{1}{2}\delta\right)&\text{otherwise}\end{cases}$
where $\delta>0$ is a tunable threshold that controls the transition between the two regimes. We search over $\delta\in\{0.01,0.05,0.1,0.25,0.5\}$ to select the value that yields the best fit.
Additionally, we introduce floor and ceiling parameters to bound the predicted values and model saturation behavior. The floor accounts for the irreducible failure rate even at high $K$, while the ceiling models the upper bound on achievable performance (e.g., due to imperfect verifiers or ambiguous problems). The final fitted form is:
[…]
We can use an unbiased estimator to evaluate best-of-$k$ selection accuracy when a fixed verifier is used to rank responses, as described in [75]. However, in the case of Weaver, the development set constitutes 1% of the data and is itself selected based on the value of $K$. In turn, the ranking of responses is no longer independent of $K$, introducing bias into the best-of-$k$ estimate.As a result, we instead rely on Monte Carlo estimates to approximate best-of-$k$ performance, sampling $k$ responses multiple times and computing the average accuracy of the top-ranked output under the $K$-dependent verifier. We use an unbiased estimator for coverage, as described in [13].
Fig. 16 and Fig. 17 along with Table 18 illustrate how the different verification strategies scale with the number of generations and the fit to the parametric form in Eq. 14. Each method exhibits characteristic scaling behavior that aligns with Eq. 14. Weaver demonstrates improved performance over naive ensembles and majority voting. The fitted parameters in Table 18 quantitatively capture these trends across datasets, providing evidence that the parametric from in Eq. 15 closely model empirical outcomes. Fig. 18 and Fig. 19 along with Table 19 illustrate the predictive performance of the parametric form in Eq. 15, showing that models fit on subsets of $K$ can extrapolate to unseen values of $K$.
Figure 16: Weaver Scaling trend fit for 70B models
Figure 17: Weaver Scaling trend fit for 8B models.
Table 18: Fitted parameters for Scaling Trends in Fig. 16 and Fig. 17.
Dataset Approach Equation floor ceil $\zeta$ $\alpha$ $\pi$ $\gamma$ R2 fit MSE fit $\delta$
GPQA-v2-Diamond (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.0000 0.9429 0.7603 0.3475 X X 0.9999 0.0000 0.5000
GPQA-v2-Diamond (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.3958 0.6728 0.7320 1.5865 0.3250 0.5053 0.9994 0.0000 0.1000
GPQA-v2-Diamond (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4283 0.4710 0.0499 1.0000 0.1217 1.0091 0.8634 0.0000 0.0100
GPQA-v2-Diamond (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.3921 0.6071 0.6553 1.9147 0.4224 0.5000 0.9975 0.0000 0.2500
MATH-500-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.6262 1.0000 0.8394 0.6427 X X 0.9994 0.0000 0.2500
MATH-500-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7870 0.9371 3.3908 3.0000 0.2869 0.5000 0.9958 0.0000 0.1000
MATH-500-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7747 0.8238 10.0000 2.1433 0.0885 2.4951 0.8655 0.0001 0.1000
MATH-500-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7883 0.9282 4.0033 3.0000 0.2573 0.5000 0.9961 0.0000 0.0100
MMLU-Pro-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.0000 0.9828 0.3303 0.3465 X X 0.9967 0.0000 0.2500
MMLU-Pro-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6912 0.9148 1.5284 3.0000 0.2764 0.5000 0.9987 0.0000 0.1000
MMLU-Pro-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6933 0.7399 0.0498 1.0001 0.1531 1.0123 0.9451 0.0000 0.0100
MMLU-Pro-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6969 0.8874 1.7834 3.0000 0.2403 0.5000 0.9944 0.0000 0.2500
MMLU-College-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.5924 0.9744 0.5071 0.5682 X X 0.9982 0.0000 0.0100
MMLU-College-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8234 0.9477 4.4129 3.0000 0.3622 0.5000 0.9987 0.0000 0.0100
MMLU-College-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8197 0.8412 0.0498 1.0001 0.2057 1.0173 0.8912 0.0000 0.0100
MMLU-College-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8235 0.9266 3.8766 3.0000 0.3012 0.5000 0.9925 0.0000 0.0500
GPQA-v2-Diamond (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.2262 0.9926 2.5454 0.8474 X X 0.9996 0.0000 0.1000
GPQA-v2-Diamond (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2463 0.4549 0.0534 1.0020 0.1953 0.7089 0.9948 0.0000 0.0500
GPQA-v2-Diamond (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2783 0.3029 1.2431 0.6756 0.0630 2.5000 0.5929 0.0000 0.0500
GPQA-v2-Diamond (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2359 0.3727 0.0701 1.0016 0.3871 0.5000 0.9408 0.0001 0.0100
MATH-500-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.4099 1.0000 1.7182 0.8949 X X 0.9976 0.0001 0.0100
MATH-500-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4110 0.7440 0.0785 1.0033 0.3333 0.5036 0.9984 0.0000 0.0100
MATH-500-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.5058 0.7038 5.1187 1.0175 0.0666 2.5000 0.9964 0.0000 0.0500
MATH-500-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.5071 0.7508 1.8068 3.0000 0.2890 0.5000 0.9796 0.0001 0.0100
MMLU-Pro-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.3906 1.0000 1.9045 0.7590 X X 0.9991 0.0000 0.0100
MMLU-Pro-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4764 0.6846 2.7876 3.0000 0.2136 0.6141 0.9985 0.0000 0.2500
MMLU-Pro-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4439 0.5662 0.1136 0.9916 0.2787 0.6659 0.9084 0.0001 0.0100
MMLU-Pro-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4771 0.7025 2.8863 3.0000 0.1728 0.5181 0.9986 0.0000 0.1000
MMLU-College-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.4316 0.9924 0.9887 0.9123 X X 0.9994 0.0000 0.5000
MMLU-College-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6226 0.8494 1.2646 3.0000 0.3346 0.5000 0.9958 0.0000 0.2500
MMLU-College-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6359 0.7368 1.0929 0.5130 0.0576 2.5000 0.9949 0.0000 0.0500
MMLU-College-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6283 0.8085 1.4279 3.0000 0.3914 0.5000 0.9845 0.0000 0.0100
Figure 18: Weaver Scaling trend predicted for 70B models
Figure 19: Weaver Scaling trend predicted for 8B models
Table 19: Fitted parameters for Scaling Trends with 90% of data in Fig. 18 and Fig. 19.
Dataset Approach Equation floor ceil $\zeta$ $\alpha$ $\pi$ $\gamma$ R2 fit MSE fit MSE pred $\delta$
GPQA-v2-Diamond (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.0000 0.9357 0.7534 0.3537 X X 0.9999 0.0000 0.0000 0.1000
GPQA-v2-Diamond (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4050 0.6756 0.9195 1.6227 0.3163 0.5000 0.9993 0.0000 0.0000 0.0500
GPQA-v2-Diamond (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4320 0.4678 0.0707 0.9864 0.0414 1.6541 0.8601 0.0000 0.0001 0.0100
GPQA-v2-Diamond (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.3809 0.6061 0.5211 1.9075 0.4360 0.5000 0.9972 0.0000 0.0000 0.0500
MATH-500-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.6639 0.9952 0.9836 0.6936 X X 0.9995 0.0000 0.0000 0.0100
MATH-500-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7869 0.9339 3.5000 3.0000 0.2985 0.5000 0.9954 0.0000 0.0000 0.0500
MATH-500-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7747 0.8241 10.0000 2.1272 0.0888 2.4999 0.8560 0.0001 0.0000 0.0500
MATH-500-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.7879 0.9221 4.2081 3.0000 0.2789 0.5000 0.9966 0.0000 0.0000 0.1000
MMLU-Pro-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.0000 0.9785 0.3263 0.3543 X X 0.9959 0.0000 0.0000 0.1000
MMLU-Pro-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6912 0.9148 1.5277 3.0000 0.2765 0.5000 0.9984 0.0000 0.0000 0.2500
MMLU-Pro-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6923 0.7379 0.0495 1.0001 0.1711 1.0148 0.9481 0.0000 0.0000 0.0500
MMLU-Pro-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6979 0.8907 1.8452 3.0000 0.2327 0.5000 0.9931 0.0000 0.0000 0.1000
MMLU-College-v2 (70B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.7723 0.9655 1.3237 0.7746 X X 0.9987 0.0000 0.0000 0.1000
MMLU-College-v2 (70B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8184 0.9436 2.2897 2.0761 0.4148 0.5000 0.9979 0.0000 0.0000 0.2500
MMLU-College-v2 (70B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8196 0.8413 0.0498 1.0001 0.2051 1.0153 0.8814 0.0000 0.0000 0.0100
MMLU-College-v2 (70B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.8234 0.9262 3.9001 3.0000 0.3036 0.5000 0.9909 0.0000 0.0000 0.1000
GPQA-v2-Diamond (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.2223 0.9976 2.4975 0.8328 X X 0.9996 0.0000 0.0000 0.2500
GPQA-v2-Diamond (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2316 0.4678 0.0559 1.0027 0.2347 0.5940 0.9965 0.0000 0.0002 0.0100
GPQA-v2-Diamond (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2773 0.2979 0.1335 0.9633 0.0479 2.5000 0.5476 0.0000 0.0001 0.0500
GPQA-v2-Diamond (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.2871 0.3845 8.5633 3.0000 0.2431 0.5000 0.9643 0.0000 0.0001 0.0100
MATH-500-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.3986 1.0000 1.6393 0.8786 X X 0.9976 0.0001 0.0001 0.0100
MATH-500-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4149 0.7417 0.0750 1.0042 0.3261 0.5179 0.9981 0.0000 0.0000 0.0100
MATH-500-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.5061 0.6976 5.9901 1.1205 0.0782 2.5000 0.9960 0.0000 0.0000 0.0500
MATH-500-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.5009 0.7342 1.6358 3.0000 0.3321 0.5000 0.9803 0.0001 0.0005 0.0100
MMLU-Pro-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.3877 1.0000 1.8798 0.7549 X X 0.9990 0.0000 0.0000 0.1000
MMLU-Pro-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4770 0.6937 3.1648 3.0000 0.2172 0.5705 0.9986 0.0000 0.0001 0.0500
MMLU-Pro-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4663 1.0000 2.4923 0.1016 0.0362 2.5000 0.9600 0.0001 0.0002 0.0500
MMLU-Pro-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.4777 0.7207 3.0170 3.0000 0.1613 0.5000 0.9987 0.0000 0.0000 0.0100
MMLU-College-v2 (8B) Pass@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})$ 0.4442 0.9912 1.0258 0.9265 X X 0.9993 0.0000 0.0000 0.5000
MMLU-College-v2 (8B) Weaver $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6178 0.8447 1.1473 3.0000 0.3523 0.5000 0.9957 0.0000 0.0001 0.2500
MMLU-College-v2 (8B) Majority1@K $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6254 0.7186 0.5765 0.9281 0.2058 1.2287 0.9736 0.0000 0.0001 0.0100
MMLU-College-v2 (8B) Naive Ensemble $y=\mathrm{floor}+(\mathrm{ceil}-\mathrm{floor})\cdot\exp(-\zeta\cdot K^{-\alpha})(1-(1-\pi)^{K^{\gamma}})$ 0.6074 0.9813 1.2560 0.1636 0.3060 2.4651 0.9988 0.0000 0.0000 0.0100
C.4 Scaling Candidate Generations
Figure 20: False Positive Rates across Verification Systems
Table 20: Weaver with 8B Models Exceeds Majority Voting and Naive Ensemble across All Datasets: Candidate are generated with Llama 3.1 8B Instruct while the weak verifiers are 8B parameters or smaller in size.
Methodology Generations ($K$) Datasets Average
MATH
500
GPQA
MMLU
College
MMLU
Pro
Baselines First Sample 1 49.8% 28.3% 64.1% 46.6% 47.2%
Majority Voting 100 69.0% 30.5% 72.7% 56.4% 57.2%
Top-Ranked RM from RewardBench [44] 100 73.8% 25.4% 70.1% 53.4% 55.7%
Top-10 RM Ensemble from RewardBench [44] 100 70.2% 22.1% 73.9% 49.4% 53.9%
Multi-Agent Verification [46] 100 65.4% 31.4% 70.5% 55.2% 55.6%
Self-Verification [99] 100 71.4% 32.2% 70.4% 53.0% 56.8%
Weaver
100 80.0% 47.1% 85.7% 67.2% 70.0%
GPT-4o-mini 1 76.8% 38.4% 82.2% 61.8% 64.8%
Claude 3.5 Haiku 1 70.0% 36.4% 75.9% 65.2% 61.9%
Oracle Verifier (Pass@100) 100 99.2% 95.2% 98.5% 96.8% 97.4%
Figure 21: Weaver Scaling - 8B Generations and Models
C.5 Scaling Verifier Count
In Table 21, we include results of scaling verifier scores. We note that for reward models (RMs), which are typically deterministic [44, 77], multiple scores must be obtained by varying the prompt; for LM Judges, we can vary either the prompt or the sampling temperature to generate diverse outputs from the same model (Table 21). We find that for both types of weak verifiers, RMs and LM judges, scaling the number of models yields better performance than sampling multiple evaluations from the same model via prompt tuning or temperature variation. However, we note that these approaches are complementary.
When breaking down weak verifiers into RMs or LM Judges, individually, we find that additional LMs leads average gains of 5.4% and 6.1%, respectively (Table 21). In contrast, sampling additional scores from a single RM or LM judge yields only 0.8% and 1.1% gains on average. These results suggest that leveraging the complementary strengths of multiple verifiers can be more effective than eliciting multiple judgments from a single verifier. Sec. C.7 provides additional details on the verifier prompting. Finally, Figure 22 illustrates the tradeoff of scaling the number of verifiers versus increasing the number of scores from a single verifier, showing that scaling verifiers is helpful when the coverage increases as we increase sample count.
Table 21: Ensembling with Multiple Verifiers Outperforms Increased Sampling with Single Verifier: Candidate responses are generated with Llama 3.3 70B Instruct while the weak verifiers range in size from 8B to 72B parameters. For details on prompting, please see Appendix C.7.
Methodology Benchmarks
MATH500 GPQA MMLU Pro
First Sample 78.0% 42.9% 69.9%
Majority Voting 83.0% 47.4% 74.4%
Best Reward Model
(1 Score)
94.4% 58.4% 81.8%
Best Reward Model
(5 Scores, 5 Prompts)
93.2% 55.3% 82.5%
Top-5 Most Accurate
Reward Models
95.4% 64.1% 87.3%
Best LM Judge
(1 Score)
90.2% 61.1% 79.5%
Best LM Judge
(5 Scores, 5 Prompts)
88.1% 57.2% 80.8%
Top-5 Most Accurate
LM Judges
93.4% 65.2% 85.4%
Figure 22: Weaver Performance Improvements from Scaling Generations and Verifiers: Increased candidate generations and weak verifiers available generally improves performance.
In Fig. 22 illustrates how the number of verifiers and repeated generations interact to influence success rate. We observe that increasing the number of generations tends to be more effective than increasing the number of verifiers alone—but only when paired with the right verification strategy. For example, naive ensembling of verifiers plateaus in performance even as more generations are added, whereas Weaver continues to improve with both axes. This highlights that generation diversity is a stronger driver of performance than verifier count alone, and that weak supervision methods like Weaver are essential to fully leverage this diversity. We illustrate the verification generation tradeoff for additional datasets in Appendix C.3.
C.6 Weaver Distillation
Figure 23: Overview of Weaver Distillation (Section 6)
For the loss function in Weaver distillation, we utilized cross-entropy loss with Adam [40]. Our classification architecture comprises a single linear classification layer with 0.1 dropout applied to the input, which consists of the final hidden state from the $[CLS]$ token. Regarding learning dynamics, we implemented linear warmup and linear decay via the Sentence-Transformers library [65], employing a learning rate of 5e-6 and training batch size of 64 across all experimental setups.
Figure 24: Weaver Distilled - Pareto Frontiers: We train/evaluate on an 80:20 split.
Table 22: Distillation Comparison of Weaver and Naive Ensemble Across Different Training Set Sizes
Methodology Dataset Training Set as Percentage of Entire Dataset Full System
5% 10% 20% 50% 80%
Weaver MATH500 78.4% 80.7% 83.9% 88.2% 91.4% 93.4%
GPQA Diamond 42.6% 46.8% 52.7% 63.1% 71.8% 73.2%
MMLU College 83.5% 85.2% 87.6% 91.0% 93.1% 94.9%
MMLU Pro 69.2% 72.5% 76.8% 83.7% 87.8% 90.2%
NaiveEnsemble MATH500 77.8% 79.6% 82.1% 86.4% 89.1% 92.4%
GPQA Diamond 42.1% 44.7% 48.9% 56.2% 62.8% 66.2%
MMLU College 84.0% 85.3% 87.2% 90.8% 93.5% 95.1%
MMLU Pro 69.5% 71.8% 74.9% 80.3% 84.7% 87.4%
C.7 Individual Verifier Optimization
While Weaver primarily focuses on aggregating multiple weak verifiers to improve overall verification quality, this appendix explores complementary techniques for optimizing individual verifiers. As mentioned earlier in the paper, existing weak verifiers often suffer from high false positive rates [79], which can limit their effectiveness, even within an ensemble.
As we scale the number of repeated samples and employ multiple verifiers, the precision of each individual verifier becomes increasingly important relative to recall. When many candidate solutions are available, a verifier can afford to miss some correct solutions (false negatives) as long as its positive predictions are highly reliable (high precision).
This observation motivates exploring methods to enhance individual verifier quality through methods such as prompt optimization — tailoring verifier prompts to maximize performance, particularly precision, with minimal or no labeled data.
C.7.1 LM Judge Prompt Optimization
LM judges often suffer from biases such as position bias (favoring answers in certain positions), verbosity bias (preferring longer answers), and self-enhancement bias (preferring answers similar to their own generation patterns) [101, 45], suggesting sensitivity to system and input prompt design.
Throughout our Weaver experiments, we used fixed, manually engineered prompts for our LM judge verifiers. However, optimizing these prompts could potentially improve individual verifier precision and reliability. Multi-Agent Verification [46] demonstrates this by crafting specialized prompts for specific verification aspects.
We explored systematically optimizing verifier prompts using DSPy [39], an open-source library that provides algorithms for optimizing language model prompts through discrete search over prompt candidates guided by a metric function. DSPy optimization works by generating, evaluating, and refining prompts that maximize task performance on a small labeled dataset.
Experimental Setup: We investigate two dimensions of prompt optimization: (1) optimization space scaling, where we progressively expand what the optimizer can modify from system instruction only (0-shot) to including 3 demonstrations (3-shot) and 5 demonstrations (5-shot); and (2) training data size scaling, where we vary labeled data from 1% to 16% to determine how much data is necessary for effective prompt optimization.
[…]
Figure 25: LM judge prompt optimization using 250 labeled examples consistently yields precision gains. Baseline methods (CoT and Custom) are compared against DSPy-optimized prompts with varying numbers of demonstrations (0-shot, 3-shot, and 5-shot).
Results: Figure 25 shows results across different datasets and optimization configurations. While we don’t observe clear scaling relationships across all datasets (possibly due to the increased stochasticity of LLM-based optimization), we observe an average precision gain of 3.8% of the best judge over the chain-of-thought (CoT) baseline judge. MATH500 shows the largest jump in precision of 9% and shows clear improvement in precision as the optimization space is scaled.
Figure 26: Scaling LM judge prompt optimization training data leads to modest precision gains. The x-axis shows the percentage of training data used (log scale), and the y-axis shows precision.
The scaling behavior with training data size (Figure 26) shows slight log-linear improvements in precision as we increase training data, though gains differ by dataset. MMLU-College shows minimal benefit from additional data, while the remaining datasets see an average boost of 3.2% in precision when scaling the training data size from 1% to 16% of the original dataset.
Figure 27: Optimized prompts often improve LM judge performance by reducing false positive rates.
(Figure 27) reveals that optimized prompts often improve both precision and accuracy by reducing false positive rates - essentially making judges more conservative in their correctness assessments. This is particularly valuable in the repeated sampling regime, where higher precision improves overall verification quality.
These findings suggest that prompt optimization can be a valuable complement to Weaver’s aggregation approach. Even with limited labeled data, targeted prompt engineering can enhance individual verifier quality, benefiting the ensemble as a whole. Further research is needed to define a more systematic recipe for verifier prompt optimization. Additionally, it remains a question of whether we can extend prompt optimization to discriminative reward models to enjoy similar gains in performance.
Appendix D Miscellaneous
D.1 Compute Requirements
Hardware Infrastructure. Our experiments were conducted using 4 compute nodes, each equipped with 8 NVIDIA H100 GPUs (80GB HBM3 memory per GPU), for a total of 32 H100 GPUs. Each node was configured with high-bandwidth NVLink connections between GPUs and inter-node communication was facilitated via NVIDIA NVLink Switch System to minimize communication overhead during distributed training and inference.
Model Parallelism and Distribution. For our 72B parameter language models, we employed a hybrid parallelism strategy combining tensor parallelism, pipeline parallelism, and data parallelism:
  • 8-way tensor parallelism across GPUs within each node
  • 4-way pipeline parallelism across nodes
  • Data parallelism for batch processing
Storage Requirements. Processing datasets of 100GB+ required significant storage infrastructure:
  • 4TB NVMe SSDs per node for dataset caching and checkpoints
  • 100TB shared network storage for full dataset repository
Software Stack. Our experiments were powered by:
  • NVIDIA CUDA 12.2
  • PyTorch 2.1 with NVIDIA NCCL for distributed communication
  • DeepSpeed ZeRO Stage 3 for memory optimization
  • Distributed data loading with webdataset format for efficient streaming